### v1

# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

Abstract
This paper focuses on the problem of automatic link selection in multi-channel multiple access control using bandit feedback. In particular, a controller assigns multiple users to multiple channels in a time slotted system, where in each time slot at most one user can be assigned to a given channel and at most one channel can be assigned to a given user. Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed probability $f_{i,j}$. The failure probabilities are not known to the controller. The assignments are made dynamically using success/failure feedback. The goal is to maximize the time average utility, where we consider an arbitrary (possibly nonsmooth) concave and entrywise nondecreasing utility function. The problem of merely maximizing the total throughput has a solution of always assigning the same user-channel pairs and can be unfair to certain users, particularly when the number of channels is less than the number of users. Instead, our scheme allows various types of fairness, such as proportional fairness, maximizing the minimum, or combinations of these by defining the appropriate utility function. We propose two algorithms for this task. The first algorithm is adaptive and gets within $\mathcal{O}(\log(T)/T^{1/3})$ of optimality over any interval of $T$ consecutive slots over which the success probabilities do not change. The second algorithm has faster $\mathcal{O}(\sqrt{\log(T)/T})$ performance over the first $T$ slots, but does not adapt well if probabilities change.
Index Terms:
Multi-armed bandit learning; Proportional fairness; Network utility maximization; Optimization; Stochastic control; Adaptive learning
I Introduction
We consider the Multiple Access Control (MAC) problem with $n$ users and $m$ channels in slotted time $t\in\mathbb{N}$. In each time slot, a controller has to assign the users to channels such that at most one user is assigned to given channel and at most one channel is assigned to a given user. The channel assignments may fail. In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the assignment fails independently with probability $1-q_{i,j}$. The controller does not know the probabilities $q_{i,j}$. Instead, at the end of every slot, it receives feedback on whether the transmission for each assigned user-channel pair succeeded or failed.
Define the matrices $\boldsymbol{Y}(t),\boldsymbol{S}(t)\in\{0,1\}^{n\times m}$ and vector $\boldsymbol{X}(t)\in\{0,1\}^{n}$, where
$S_{i,j}(t)=\begin{cases}1&\text{ if link }i,j\text{ is successful in time slot }t\\ 0&\text{ otherwise},\end{cases}$
$Y_{i,j}(t)=\begin{cases}1&\text{ user }i\text{ is assigned to channel }j\text{ in time slot }t\\ 0&\text{ otherwise},\end{cases}$
and $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for all $i\in\{1,2,\dots,n\}$. The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the link failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.  [^1] The limit is assumed to exist for simplicity of this introduction; the precise goal is to maximize a $\lim\inf_{T\to\infty}\phi(\boldsymbol{\overline{X}}(T))$
We also focus on establishing finite time bounds. In particular, for given a finite time horizon $T\in\mathbb{N}$, we require the algorithm to satisfy
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)\leq g(T),$
where $\phi^{\text{opt}}$ is the optimal utility of the original problem and $g$ is a nonnegative function such that
$\lim_{T\to\infty}g(T)=0$
In addition, we are looking for algorithms that are adaptive. Formally, consider a system in which the channel success probabilities may change. In such a system, given a $T\in\mathbb{N}$, we require
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$
for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $[T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame. Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $[T_{0}:T_{0}+T-1]$ of the above scenario. Note that $g$ is the same function regardless of $T_{0}$.
I-A Utility functions
One utility function that can be used in our work is $\phi(\boldsymbol{x})=\min\{x_{1},\dots,x_{n}\}$. This is a nonsmooth utility function that seeks to maximize the minimum time-average success rate across all users. However, if there is one user with very low success probability, this utility function can cause almost all the resources to be devoted to that user, resulting in poor performance for all users.
Another choice is $\phi(\boldsymbol{x})=w_{1}\min\{x_{1},\dots,x_{n}\}+w_{2}\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $w_{1},w_{2},\beta$ are given nonnegative weights. The logarithmic term introduces a form of proportional fairness [1, 2]. See also discussion of different utility functions in [1, 2, 3, 4, 5]
I-B Related work
Network scheduling in stochastic environments is a widely considered problem in the literature on communication networks. Some examples include scheduling in wireless networks [6, 7], computer networks [8], vehicular networks [9] and unmanned aerial vehicle networks [10]. These works consider different goals such as optimal scheduling for queue-stability [11], power minimization [12], utility maximization [13], ensuring fairness [14], flow control [15], and multiple access control [16].
Multiple access, where many users access a limited number of communication channels, is an important problem in network scheduling. Here, it is desirable for users to be scheduled to avoid collisions. Link failures occur when the receiver is unable to decode the packet transmissions. This can occur, for example, when a fixed transmission power is used, but channel conditions have random and unknown fluctuations so that the received signal strength is insufficient for decoding. Different links can have different properties (such as different geographic distances to the receiver), so they can also have different success probabilities. The naive approach of assigning the users with links with the least failure probabilities leads to unfairness since, in such a scenario, users with high link failure probabilities will never be assigned. A common approach to solve this problem is to maximize a utility function of the time-averaged success. This problem has been well studied in the full information scenario when either the fluctuating channel conditions are known before transmission (called opportunistic scheduling) or when channel success probabilities are known in advance. Opportunistic scheduling has been considered using utility functions [17], Lyapunov drift [18], Frank-Wolfe [19, 20, 21], primal-dual [22, 23], and drift-plus-penalty [24]. The case when success probabilities are known in advance can be solved offline as a convex optimization problem using the mirror descent technique [25].
The problem becomes challenging when the success probabilities are not known, but we only receive bandit feedback on the successes. The problem has to be approached by combining ideas from optimizing functions of time averages with multi-armed bandit learning. The work on bandits with vector rewards and concave utility functions can be adapted for the single channel case ($m=1$) of our problem (See [26, 27, 28]). There are two main drawbacks to these approaches. First, the above works do not consider the matching constraints considered in our work. Next, they focus on upper confidence bound (UCB) techniques and are not adaptive. It is possible to develop adaptive algorithms using techniques from the above works together with the EXP3 algorithm [29] for the single channel case. However, they cannot be directly extended for the multi-channel case. The main reason for this is the complexity of the inner problem arising in each iteration which is a problem over the set of doubly stochastic matrices (Birkhoff polytope). This can be addressed using Sinkhorn’s algorithm [30] in each iteration. However, this results in an algorithm that requires computationally complex inner iterations to be performed within each iteration. The work of [31] uses follow the regularized leader approach to solve adversarial bandit problems over the set of doubly stochastic matrices. However, their algorithm relies on a computationally expensive inner iteration similar to what we described before. The work of  [32] proposes an algorithm to solve online optimization problems over transport polytopes. All the inner iterations of their algorithm have explicit solutions thanks to the rounding trick introduced for transport polytopes by [33]. In our work, we adapt the rounding trick to the Birkhoff polytope and develop an algorithm for general $m$. We also develop a UCB-based algorithm for our problem. It should be noted that although this algorithm achieves faster convergence, it is not adaptive. However, the UCB algorithm has additional advantages such as distributed implementation. Our paper also treats general (possibly nonsmooth) utility functions, so Frank-Wolfe methods that rely on smoothness cannot be used.
Existing work on using learning for scheduling problems in communications include spectrum sharing [34, 35], network utility maximization with unknown utility functions [36], and queue stability [11]. The existing literature on bandit learning-based methods for scheduling problems focuses on maximizing throughput [37]. The work of [36] considers bandit learning with unknown utilities, and the work of [38] considers the online resource allocation problem, where the objective is to maximize the $\alpha$-fairness of the allocations. Neither of the above works considers the bandit reward structure, and both works have a different problem structure from ours. In addition, the above works do not focus on the adaptiveness of the proposed algorithms.
It should be noted that the concept of adaptiveness considered in this paper is different from the adversarial setting [39, 40, 29], although they have similarities. In adversarial settings, optimality is defined with respect to the rewards received throughout the time horizon. This is useful when no stochastic assumptions can be placed on the rewards. On the other hand, the adaptiveness considered in this paper is useful when we can place stochastic assumptions on the rewards, but the distributions may change from time to time. In this case, we can define an optimal strategy for each time frame in which the reward distribution remained constant. Hence, the goal in this frame is to learn the aforementioned strategy irrespective of the reward distributions of the past. An approach commonly used in problems of this flavor is minimizing dynamic regret [41, 42, 43]. Here, the regret is modified to account for the changing environments and the regret bounds are in terms of some measure that captures the degree of change. Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [41, 42], and sliding window-based algorithms [43]. We utilize a simpler notion of adaptiveness and develop an algorithm for the case with general utility functions.
Multi-armed bandit (MAB) learning [44, 45] is a class of extensively studied problems. In the classical MAB setting, a user chooses from multiple arms, each incurring a fixed mean cost. The choice is made without knowledge of the mean costs, using the feedback received on the cost of the chosen arm. The goal is to learn the arm with the lowest mean cost as fast as possible. Typically, the algorithms implemented to solve such problems involve an exploration phase, where the user explores all arms in order to learn the mean costs, after which, in the exploitation phase, the user exploits the learned information to make good decisions [46, 47]. Upper confidence bound-based algorithms, where the algorithm maintains an upper bound on the mean cost of each arm, is a popular approach to solve these problems [47, 48]. Adversarial bandit learning [39, 40], where an adversary is allowed to assign the cost of each arm after the user selects the arm, is an extension of multi-armed bandit learning. In such problems, no stochastic assumptions are placed on the rewards. These problems are solved using algorithms such as the EXP3 algorithm [29], where in each iteration of the algorithm, a probability distribution over the space of all possible actions is found, after which the decision is randomly sampled from the distribution. There are also many extensions of the MAB problem, such as linear bandits, contextual bandits, and combinatorial bandits [47].
I-C Our Contributions
We develop and analyze two algorithms to solve the problem of automatic link selection in multi-channel multiple access, combining the ideas of multi-armed bandit learning and Lyapunov optimization. Although the classical MAB problem has been widely analyzed for maximizing linear utilities, they suffer from lack of fairness in assignments when applied to the considered problem. It is notable that our method allows either smooth or nonsmooth concave, entrywise nondecreasing utilities.
We prove that the first algorithm gets within $\mathcal{O}(T^{-1/3}\log(T))$ of optimality over any interval of $T$ consecutive time slots during which the (unknown) success probabilities do not change. If these probabilities are different before $T_{0}$, but change to new probabilities during $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, our performance guarantees for the new interval are independent of behaviors before $T_{0}$, even though the algorithm does not know the exact time $T_{0}$ of the change. Hence, the algorithm is adaptive. This is possible thanks to the importance sampling technique in [29], which we use to estimate the true link failure probabilities in each time slot. This technique uses feedback only from the previous time slot.
The second algorithm is based on upper confidence bound (UCB) techniques and gets within $\mathcal{O}(\sqrt{\log(T)/T})$ of optimality within a finite time horizon of $T$ time slots. Although this algorithm has faster convergence compared to the first algorithm, it is not adaptive due to the UCB-based estimation. However, the channel-user assignment of this algorithm has a max-weight structure and can be solved using the well-known Hungarian algorithm [49]. This is in contrast to the first algorithm, where the assignment is sampled from a distribution in each time slot, which may be less desirable in certain scenarios. Due to the max-weight decisions, the second algorithm can be implemented in a distributed setting in the absence of a centralized controller, given that all the users have access to feedback on which channels are successfully accessed in each time slot. Simulations also depict the adaptiveness of the first algorithm and the faster convergence of the second algorithm.
I-D Notation
[…]
I-E Definitions
In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also, define the sets
$\Delta_{l,\varepsilon}=\Big{\{}\boldsymbol{p}\in\mathbb{R}^{k}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\Big{\}},\text{ where }l\in\mathbb{N},$
$\mathcal{S}^{\text{row}}_{\varepsilon}=\Bigg{\{}\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}:\sum_{k=1}^{s}P_{i,k}=1,P_{i,j}\geq\varepsilon\ \forall i,j\in[s]\Bigg{\}},$
$\mathcal{S}^{\text{col}}_{\varepsilon}=\Bigg{\{}\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}:\sum_{k=1}^{s}P_{k,j}=1,P_{i,j}\geq\varepsilon\ \forall i,j\in[s]\Bigg{\}},$
$\mathcal{S}^{\text{doub}}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cap\mathcal{S}^{\text{row}}_{\varepsilon},\ \mathcal{S}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cup\mathcal{S}^{\text{row}}_{\varepsilon}.$
We also denote $\Delta_{l}=\Delta_{l,0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$. We hide the dependence on $s$ in the notation for sets for clarity.
I-F Assumptions
Before moving on to the problem, we state our main assumptions.
  • A1
    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $[0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$. Hence, $\phi$ is $\sqrt{n}B$-Lipschitz continuous. Also, let $\phi^{\text{max}}=\max_{\boldsymbol{x}\in[0,1]^{n}}|\phi(\boldsymbol{x})|$.
  • A2
    We have access to the solution of the problem $\max_{\boldsymbol{x}\in[0,1]^{n}}[\phi(\boldsymbol{x})+\sum_{i=1}^{n}c_{i}x_{i}]$, for all $\boldsymbol{c}\in\mathbb{R}^{n}_{+}$.
    Note: This assumption is valid for most separable functions $\phi$. For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem has an explicit solution.
II Problem Setup
We formalize the problem of interest below.
(P1:) $\underset{\boldsymbol{Y}(1),\boldsymbol{Y}(2),\dots}{\text{max}}\lim\inf_{T\to\infty}\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$ (1)
[…]
$\ \ \ \ \ \ \ t,\tau\in\mathbb{N}\text{ and }\tau\geq t$ (2)
$\ \ \ \ \ \boldsymbol{Y}(t),\text{ and }S_{i,j}(\tau)\text{ are independent for all }t,\tau\in\mathbb{N},$
$\ \ \ \ \ \ \ (i,j)\in[n]\times[m],\tau<t,Y_{i,j}(\tau)\neq 1$ (3)
$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{n\times m}\ \forall t\in\mathbb{N}$ (4)
$\ \ \ \ \ \sum_{i=1}^{n}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[m]$ (5)
$\ \ \ \ \ \sum_{j=1}^{m}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[n]$ (6)
$\ \ \ \ \ X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)\forall t\in\mathbb{N},\ i\in[n],$ (7)
where constraint (2) ensures transmission decisions do not know success/failures before they happen; (3) ensures we cannot use information that is never observed. Define $s=\max\{n,m\}$ and $\phi^{\text{opt}}$ as the optimal objective value of (P1).
Lemma 1
Consider the following problem.
(P2:) $\underset{\boldsymbol{P},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (8)
[…]
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (10)
$\ \ \ \ \ \sum_{j=1}^{m}q_{i,j}P_{i,j}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (11)
and let $\phi^{*}$ denote the optimal objective value of the above problem. Then we have $\phi^{*}=\phi^{\text{opt}}$.
Proof:
We first prove that $\phi^{*}\geq\phi^{\text{opt}}$. Fix $\varepsilon>0$. Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (2)-(6), such that
$\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}^{*}(t)\}\right)\geq\phi^{\text{opt}}-\varepsilon,$ (12)
where $X^{*}_{i}(t)=\sum_{j=1}^{m}Y^{*}_{i,j}(t)S_{i,j}(t)$ for all $t\in\mathbb{N}$ and $i\in[n]$. Define the matrix $\boldsymbol{\tilde{P}}^{*}\in\mathbb{R}^{s\times s}$ such that
$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if }(i,j)\in[n]\times[m]\\ 0&\text{ otherwise. }\end{cases}$
Let us define $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\boldsymbol{P}^{*}\geq\boldsymbol{\tilde{P}}^{*}$. It can be observed that such a $\boldsymbol{P}^{*}$ exists from the definition of $\boldsymbol{\tilde{P}}^{*}$ and the constraints (4)-(6). Let $\boldsymbol{\gamma}^{*}\in[0,1]^{n}$ such that $\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}$ for $i\in[n]$. It is easy see that $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ satisfy constraints (9)-(11). Also, we have
$\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}\geq\sum_{j=1}^{m}q_{i,j}\tilde{P}^{*}_{i,j}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{X^{*}_{i}(t)\}$ (13)
since (2) ensures $\mathbb{E}\{Y^{*}_{i,j}(t)S_{i,j}(t)\}=\mathbb{E}\{Y^{*}_{i,j}(t)\}q_{i,j}$. Hence,
$\phi^{*}$ $\geq_{(a)}\phi(\boldsymbol{\gamma}^{*})\geq_{(b)}\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}^{*}(t)\}\right)\geq_{(c)}\phi^{\text{opt}}-\varepsilon,$
where (a) follows since $\phi^{*}$ is the optimal objective value of (P2) and $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is feasible for (P2), (b) follows from (13) and the entrywise nondecreasing property of $\phi$ (see assumption A1) and (c) follows from (12). The above is true for all $\varepsilon>0$. Hence, we have $\phi^{*}\geq\phi^{\text{opt}}$ as desired.
Now, we prove that $\phi^{*}\leq\phi^{\text{opt}}$. Let $\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*}$ denote the optimal solution for (P2). Using Birkhoff-von Neumann Theorem [50], notice that there exists $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}^{*}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, and $\boldsymbol{s}\in\Delta_{r}$. In each time slot, we simply sample $l_{t}\sim\boldsymbol{s}(t)$ and set $\boldsymbol{Y}(t)$ according to $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for $i\in[n],j\in[m]$. It can be easily seen that this policy yields an objective value of $\phi^{*}$ for (P1). Hence, we are done. ∎
III Adaptive Algorithm
Now, we develop our first algorithm. The idea is to find $\boldsymbol{\tilde{P}}(t)$ as a column stochastic matrix and a row stochastic matrix alternatively in odd and even slots. Then we obtain $\boldsymbol{P}(t)$ by approximating $\boldsymbol{\tilde{P}}(t)$ by a doubly stochastic matrix, using the rounding trick similar to the one introduced in [33], after which we sample a permutation matrix from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [50]. To do the assignment in time slot $t-1$, we discard the last $s-n$ rows or last $s-m$ columns of the permutation matrix, depending on whether $m$ or $n$ is larger.
1 Initialize $\boldsymbol{\tilde{P}}(1)\in\mathcal{S}^{\text{doub}}_{\varepsilon}$ and the virtual queues $\boldsymbol{Q}(1)\in[0,BV+1]^{n}$ arbitrarily.
2 for each time slot $t\in\mathbb{N}$ do
3       Set $\boldsymbol{P}(t)=\text{ROUND}(\boldsymbol{\tilde{P}}(t))$ (This function yields $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ ).
4
5      Using Birkhoff-von Neumann Decomposition [50], find $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}(t)=\sum_{l=1}^{r}s_{l}(t)\boldsymbol{M}^{l}$, and $\boldsymbol{s}(t)\in\Delta_{r}$.
6       Sample $l_{t}\sim\boldsymbol{s}(t)$ and take action $\boldsymbol{Y}(t)$, where $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for all $i\in[n]$, $j\in[m]$, and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
7       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=S_{i,j}(t)Y_{i,j}(t)/P_{i,j}(t)$ for all $i\in[n]$, and $j\in[m]$.
8       Find $\boldsymbol{\gamma}(t+1)\in[0,1]^{n}$, and $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{Q}$ using
$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big{]}.$
$\boldsymbol{\tilde{P}}(t+1)=\arg\min_{\boldsymbol{P}\in\mathcal{Q}}\Big{[}-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\Big{]},$ (14)
where $\mathcal{Q}=\mathcal{S}^{\text{col}}_{\varepsilon}$ if $t$ is even, and $\mathcal{Q}=\mathcal{S}^{\text{row}}_{\varepsilon}$ if $t$ is odd, and the divergence $D$ is defined by
$D(\boldsymbol{X}\lVert\boldsymbol{Y})=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right).$ (15)
for all $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}.$
9       Update the virtual queues
$\boldsymbol{Q}(t+1)=\left[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right]_{+},$ (16)
where $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for $i\in[n]$.
10
Algorithm 1 Adaptive MAC
Below, we introduce the ROUND function, a technique adapted from the one introduced in [33] to approximate a nonnegative matrix by a matrix in the transport polytope. The technique can be readily extended to the Birkhoff polytope using the same algorithm.
ROUND($\boldsymbol{P}$) function for $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$ :
  1. 1.
    Define the matrix $\boldsymbol{P}^{{}^{\prime}}$ (row normalization of $\boldsymbol{P}$) using
    $P^{{}^{\prime}}_{i,j}=\begin{cases}\frac{P_{i,j}}{\sum_{l=1}^{s}P_{i,l}}&\text{ if }\sum_{l=1}^{s}P_{i,l}>1\\ P_{i,j}&\text{ otherwise.}\end{cases}$ (17)
  2. 2.
    Define the matrix $\boldsymbol{P}^{{}^{\prime\prime}}$ (column normalization of $\boldsymbol{P}^{{}^{\prime}}$) using
    $P^{{}^{\prime\prime}}_{i,j}=\begin{cases}\frac{P^{{}^{\prime}}_{k,j}}{\sum_{k=1}^{s}P^{{}^{\prime}}_{k,j}}&\text{ if }\sum_{k=1}^{s}P^{{}^{\prime}}_{k,j}>1\\ P^{{}^{\prime}}_{i,j}&\text{ otherwise.}\end{cases}$ (18)
  3. 3.
    Define the output matrix $\boldsymbol{Q}$,
    $\boldsymbol{Q}=\begin{cases}\boldsymbol{P}^{{}^{\prime\prime}}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}}{C}&\text{ if }C\neq 0\\ \boldsymbol{P}^{{}^{\prime\prime}}&\text{ otherwise, }\end{cases}$ (19)
    where $C=\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}$.
It can be shown that ROUND$(\boldsymbol{P})\in\mathcal{S}_{\varepsilon/s}^{\text{doub}}$ whenever $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$.
In Algorithm 1, we provide the algorithm for the task. In the following section, we focus on solving the intermediate problem (8), after which we move on to the analysis of the Algorithm. Problem (P2) motivates the definition of auxiliary variables $\boldsymbol{\gamma}(1),\boldsymbol{\gamma}(2),\dots\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(1),\boldsymbol{Q}(2),\dots$ to aid the algorithm. In particular, we will show that we can use our algorithm to get arbitrarily close to $\phi^{\text{opt}}$ in objective value.
III-A Solving Problem (8)
Finding $\boldsymbol{\gamma}(t+1)$: Notice that this problem can be solved due to the Assumption A2.
Finding $\boldsymbol{\tilde{P}}(t+1)$: We will only consider the case when $t$ is even. The case when $t$ is odd can be solved similarly. Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$. To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
$x_{i}=\begin{cases}\eta Q_{i}(t)\hat{S}_{i,j}(t)&\text{ if }i\in[n],j\in[m]\\ 0&\text{ otherwise}\end{cases}$
and $\boldsymbol{y}$ is the $j$-th column of $\boldsymbol{\tilde{P}}(t)$. The problem to be solved is
(P3:) $\underset{\boldsymbol{p}}{\text{min}}-\sum_{i=1}^{s}x_{i}p_{i}+D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{y})$
[…]
where $D_{\text{KL}}(\cdot\lVert\cdot)$ in this case is the KL-divergence. It should be noted that (P3) has a classic structure that is solved in [11]. In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$. First, assume that $\boldsymbol{z}$ is sorted in the increasing order. Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
$u^{i}_{j}=\begin{cases}\varepsilon&\text{ if }j\leq i\\ \frac{z_{j}}{\sum_{l=i+1}^{s}z_{l}}(1-\varepsilon i)&\text{ if }j>i\\ \end{cases}$
satisfies, $u^{i}_{j}\geq\varepsilon$ for all $j\in[i+1:s]$. Then it can be shown that $\boldsymbol{u}^{i}$ is the solution to (P3). Hence, solving (P3) amounts to calculating $\boldsymbol{u}^{i}$ for each $i\in[0:s-1]$ and checking the above condition.
We first establish the following two lemmas regarding the ROUND operation. A version of the following lemma is proved for the transport polytope in [33]. We adapt the idea for the Birkhoff polytope.
Lemma 2
The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$, intermediate matrices $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function in Algorithm 1 satisfy
  1. 1.
    $\boldsymbol{Q}\in\mathcal{S}^{\text{doub}}$.
  2. 2.
    If the input $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$, we have $\boldsymbol{Q}\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$.
  3. 3.
    $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$.
  4. 4.
    $\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right)$.
Proof:
See Appendix A. ∎
Next, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
Lemma 3
We have $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ and
$\lVert\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)\rVert_{1}\leq 2\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}.$
Proof:
The fact that $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ follows directly from Lemma 2-2.
For the other part, we only consider the case when $t$ is even. The case when $t$ is odd follows similarly.
Claim 1: When $t$ is even, we have
$\lVert\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)\rVert_{1}\leq 2\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}.$
This again follows by direct application of Lemma 2-4. (Notice that $\lVert\boldsymbol{\tilde{P}}(t)\boldsymbol{1}-\boldsymbol{1}\rVert_{1}=0$, since $\boldsymbol{\tilde{P}}(t)\in\mathcal{S}^{\text{row}}$ when $t$ is even)
Claim 2: When $t$ is even, we have
$\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\leq\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}.$
Notice that
$\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}|\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)|\geq_{(a)}\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}\tilde{P}_{i,j}(t+1)-\sum_{i=1}^{s}\tilde{P}_{i,j}(t)\right\rvert$
$=\lVert\boldsymbol{\tilde{P}}(t+1)^{\top}\boldsymbol{1}-\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}\rVert_{1}=\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1},$
where (a) follows from the triangle inequality and the last equality follows since $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{S}^{\text{col}}$ when $t$ is even.
Combining the two claims, we are done. ∎
Now, we introduce some useful preliminary lemmas. First, we introduce the push-back lemma regarding minimizing strongly convex functions (See, for example [51]).
Lemma 4 (Push-back)
Consider $\mathcal{Q}\in\{\mathcal{S}^{\text{doub}}_{\varepsilon},\mathcal{S}^{\text{row}}_{\varepsilon},\mathcal{S}^{\text{col}}_{\varepsilon},\mathcal{S}\}$, and $\varepsilon\in[0,1/s]$. Let $g:\mathbb{R}_{+}^{s\times s}\to\mathbb{R}$ be a convex function. Fix $\alpha>0$, and $\boldsymbol{Y}\in\mathcal{Q}$. Let
$\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}\left[g(\boldsymbol{X})+\alpha D(\boldsymbol{X}\lVert\boldsymbol{Y})\right],$
where $D(\cdot\lVert\cdot)$ is the divergence defined in (15). Then,
$g(\boldsymbol{X}^{*})+\alpha D(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}),$
for all $\boldsymbol{Z}\in\mathcal{Q}$.
We now establish the following two lemmas.
Lemma 5
We have for $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}$,
$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert^{2},$ (20)
where $D(\boldsymbol{X}\lVert\boldsymbol{Y})$ is the divergence defined in (15).
Proof:
Notice that
$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\sum_{i=1}^{s}\frac{1}{2}\left(\sum_{j=1}^{s}|X_{i,j}-Y_{i,j}|\right)^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}$
where the first inequality follows from the Pinsker’s inequality and the last inequality follows from the fact that for $a_{1},a_{2},\dots,a_{s}\in\mathbb{R}$, $\sum_{i=1}^{s}a_{i}^{2}\geq\left(\sum_{i=1}^{s}a_{i}\right)^{2}/s$. ∎
Next, we state a result on the properties of divergence D.
Lemma 6
We have $D(\boldsymbol{X}\lVert\boldsymbol{Y})\leq s\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{X}\in\mathcal{S}$, where $\boldsymbol{Y}\in\mathcal{S}_{\varepsilon}$.
Proof:
Notice that
$D(\boldsymbol{X}\lVert\boldsymbol{Y})$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right)\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{1}{\varepsilon}\right)=s\log\left(\frac{1}{\varepsilon}\right),$
where (a) follows since $X_{i,j}\leq 1$, $Y_{i,j}\geq\varepsilon$, and log is a non-decreasing function. ∎
Before moving on to the main theorem, we establish a deterministic bound on the queue size $\lVert\boldsymbol{Q}(t)\rVert$ for our Algorithm (Algorithm 1).
Lemma 7
We have for all $t\in\mathbb{N}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$, where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1 and $V$ is the utility parameter of Algorithm 1.
Proof:
We first prove the following claim.
Claim: If $Q_{k}(t)>BV$, for $k\in[n]$, then $\gamma_{k}(t+1)=0$.
Proof:
Assume the contrary $Q_{k}(t)>BV$, and $\gamma_{k}(t+1)>0$.
Consider $\boldsymbol{\tilde{\gamma}}$ such that
$\tilde{\gamma}_{i}=\begin{cases}0&\text{ if }i=k\\ \gamma_{i}(t+1)&\text{ otherwise.}\end{cases}$
Since, $Q_{k}(t)>BV$, we have
$[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (21)
due to the bounded subgradient property (Assumption A1). Hence, notice that
$\sum_{i=1}^{n}$ $[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})=\gamma_{k}(t+1)[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (22)
where the last inequality follows due to $\gamma_{k}(t+1)>0$ from assumption and (21).
The subgradient inequality for the convex function $f:[0,1]^{n}\to\mathbb{R}$ given by, $f(\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}$ gives,
$\sum_{i=1}^{n}[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})\leq f(\boldsymbol{\gamma}(t+1))-f(\boldsymbol{\tilde{\gamma}})\leq 0,$
where the last inequality follows from the optimality of $\boldsymbol{\gamma}(t+1)$ for (P3). This is a contradiction with (22). ∎
Now, we move on to the proof of Lemma 7. Fix $i\in[n]$. We use induction to prove that $Q_{i}(t)\leq BV+1$ for all $t\in\mathbb{N}$. Notice that the result trivially follows for $t=1$. Assume $Q_{i}(t)\leq BV+1$. We establish $Q_{i}(t+1)\leq BV+1$. We consider two cases.
Case 1: $Q_{i}(t)>BV$: Notice that from the previous claim, we should have $\gamma_{i}(t+1)=0$. Hence, we have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)=0$, and $X_{i}(t)\geq 0$, and the last inequality follows from induction hypothesis.
Case 2: $Q_{i}(t)\leq BV$: We have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)+1\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)\leq 1$, and $X_{i}(t)\geq 0$, and the last inequality follows from the description of the case.
Hence, we are done. ∎
Define the drift $\Delta(t)$ as
$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
and the history $\mathcal{H}(t)$ as the sigma algebra generated by
[…]
Notice that $\boldsymbol{\tilde{P}}(\tau),\boldsymbol{P}(\tau),\boldsymbol{\gamma}(\tau),\boldsymbol{Q}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable.
Now, we have the following bound on $\Delta(t)$.
Lemma 8
We have for all $t\in\mathbb{N}$,
$\Delta(t)$ $\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
Proof:
Notice that from the queuing equation (16), we have for all $i\in[n]$,
$Q^{2}_{i}(t+1)$ $\leq\left(Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\right)^{2}$
$\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t+1)+X^{2}_{i}(t)+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right]$
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right],$
where for the last inequality, we have used $X_{i}(t)\in\{0,1\}$. Summing the above for $i\in[n]$, we have
$\lVert\boldsymbol{Q}(t+1)\rVert^{2}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right].$
Taking the expectations conditioned $\mathcal{H}(t)$:
$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\mathbb{E}\{\gamma_{i}(t+1)|\mathcal{H}(t)\}-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right].$
Taking expectations, and performing simple algebraic manipulations, we have the result. ∎
Lemma 9
We have
$-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$ $\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
$\ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
where $\hat{S}_{i,j}(t)$ is defined in line 6 of Algorithm 1.
Proof:
Notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (24)
Now, we handle each of the above four terms separately. Notice that
$\sum_{i=1}^{n}$ $\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}$
$\leq_{(a)}\frac{3\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{6s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}\right\}$
$\leq\frac{3\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
where for (a) we use $\frac{3\eta s}{2}\lVert\boldsymbol{a}\rVert^{2}+\frac{1}{6\eta s}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{k}a_{i}b_{i}$ for $k$-dimensional vectors $\boldsymbol{a,b}$, and the last inequality follows from Lemma 5. Next, Notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}$$\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{12s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}$
$\leq_{(a)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}$
$\leq_{(b)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
where (a) follows from Lemma 3, and (b) follows from Lemma 5. Next, notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$ $=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
Now, using the above in (9) we have
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\}$
$\leq\frac{9\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}.$ (25)
But notice that
$\mathbb{E}$$\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i,j}(t)}{P^{2}_{i,j}(t)}Y_{i,j}(t)\right\}$
$\leq_{(a)}\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}Y_{i,j}(t)\right\}=\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(b)}\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}$
$\leq\frac{nms}{\varepsilon}(BV+1)^{2},$
where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 3, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Lemma 7. Combining with (9), we are done. ∎
Now, we introduce the following lemma that will be useful in deriving the final bounds.
Lemma 10
We have for any $T,T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
$VT\phi(\boldsymbol{\gamma})$ $-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
$\ \ \ \ +\frac{n(BV+1)^{2}}{2}.$
Proof:
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}$
$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$ (26)
where (a) follows by Lemma 9. Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (8) (see Algorithm 1) with Lemma 4, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{m}\hat{S}_{i,j}(t)P_{i,j}\right]+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))-\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1)).$
Taking expectations of the above, we have
$-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\}$
$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\gamma_{i}\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
$\ \ \ \ =-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}P_{i,j}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
Substituting the above in (10), we have
$\Delta$ $(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n+\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}.$
Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T+T_{0}-1\}$, we have
$\frac{1}{2}$ $\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(T_{0}))\right\}-\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(T+T_{0}))\right\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i,j}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right),$ (27)
where for the last inequality, we have used Lemma 6. Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Lemma 7, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Next, we have the following lemma that combines iterates $\boldsymbol{\gamma}(T_{0}+1),\boldsymbol{\gamma}(T_{0}+2)\dots,\boldsymbol{\gamma}(T+T_{0})$, and $\boldsymbol{X}(T_{0}),\boldsymbol{X}(T_{0}+1)$ $\dots,\boldsymbol{X}(T+T_{0}-1)$.
Lemma 11
We have for any $T,T_{0}\in\mathbb{N}$,
$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T}.$
Proof:
Notice that
$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)$
$=\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right)\right)$
$\leq_{(a)}\phi\left(\frac{1}{T}\Bigg{(}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}[\boldsymbol{Q}(t+1)-\boldsymbol{Q}(t)]\right)\Bigg{)}$
$\leq_{(b)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{\boldsymbol{Q}(T+T_{0})}{T}\right)$
$\leq_{(c)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T},$
where (a) follows from the entrywise nondecreasing property of $\phi$, and the queuing equation (16), for (b) we use $\boldsymbol{Q}(T_{0})\geq 0$ and the entrywise nondecreasing property of $\phi$, and (c) follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.  [^2] Assumption A1 ensures that $\phi$ can be extended to an entrywise nondecreasing function that is $\sqrt{nB}$-Lipschitz continuous over the domain $[0,\infty)^{n}$.
Now, we are ready to establish the performance bound of the algorithm.
Theorem 1
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/s)$, $\eta,V>0$, Algorithm 1 yields
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$ $\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}$
$\ \ \ \ +\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}.$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
Proof:
Substituting $\boldsymbol{P}^{*}(1-\varepsilon s)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 10, where $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of (P2) defined in Lemma 1 and $\boldsymbol{1}$ is the all 1 matrix, we have
$VT\phi(\boldsymbol{\gamma}^{*})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left(\gamma^{*}_{i}-\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}\right)\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\sum_{j=1}^{m}\left(q_{i,j}sP^{*}_{i,j}-q_{i,j}\right)\mathbb{E}\{Q_{i}(t)\}+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
$\leq_{(a)}nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\varepsilon nmsT(BV+1)+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2},$ (28)
where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Lemma 7. Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain
$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}.$
Combining with Lemma 11, we have
$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Lemma 7. Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{\text{opt}}$. Using Jensen’s inequality, we are done. ∎
IV Discussion
IV-A Adaptiveness
A careful inspection of the proof of Theorem 1 shows that the result holds even if the success probabilities $q_{i,j}$ changed before time $T_{0}$, as long as they remained constant during $[T_{0}:T_{0}+T-1]$. Hence, the adaptiveness is satisfied.
IV-B Enforcing user fairness
Consider the case when $n\geq m$ (the number of users is at least the number of channels). Assume we fix a $\theta\in(0,1/s]$ such that each user is required to transmit at least $\theta$ fraction of the time on average on each channel and no user can transmit more than $(1-(n-m)\theta)$ fraction of the time. This enables faster adaptation to the new optimality point.
To see this, notice that for (P1), we require the additional constraints of
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y_{i,j}(t)\}\geq\theta.$
for all $i\in[n]$ and $j\in[m]$ and
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}\{Y_{i,j}(t)\}\leq 1-(n-m)\theta,$
for all $i\in[n]$. This will simply transform the constraint $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$ of (P2) to $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\theta}$. Setting $\varepsilon=\theta$ in Algorithm 1, this will allow us to use $\boldsymbol{P}^{*}$ directly in (1) in Theorem 1, instead of $(1-\varepsilon s)\boldsymbol{P}^{*}+\varepsilon\boldsymbol{1}$, which gives
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\theta V}(BV+1)^{2}+\frac{\theta nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\theta}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{nB(BV+1)}{T}.$
Now, since we assume $\theta$ to be a constant, using $\eta=\Theta(1/T)$, and $V=\Theta(\sqrt{T})$, we have
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}=O\left(\frac{1}{\sqrt{T}}\right)$
IV-C Case $m=1$
This case can be approached using the work on bandits with vector rewards and concave utility functions [26]. However, the algorithms proposed are not adaptive. It turns out that the algorithm for the general multi-channel case can be simplified for the single-channel case preserving the adaptiveness. In addition, the simplified algorithm has faster convergence. The idea of the algorithm is to find a distribution $\boldsymbol{p}(t)$ over the users in the $(t-1)$-th iteration, after which we sample the user to be assigned on the $t$-th iteration from $\boldsymbol{p}(t)$. Similar to the multi-channel case, we use auxiliary variables $\boldsymbol{\gamma}(t)\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(t)\in\mathbb{R}^{n}_{+}$ for all $t\in\mathbb{N}$. For notational convenience, for each $t\in\mathbb{N}$ and $i\in[n]$, we will use $q_{i},S_{i}(t)$ and $Y_{i}(t)$ instead of $q_{i,1},S_{i,1}(t)$ and $Y_{i,1}(t)$, respectively. The algorithm uses three parameters $\varepsilon\in(0,1/n)$, $V>0$, and $\eta>0$. See Algorithm 2 for details of the implementation.
1 Initialize $\boldsymbol{p}(1)\in\Delta_{n,\varepsilon}$, $\boldsymbol{\gamma}(1)\in[0,1]^{n}$, and the virtual queues $\boldsymbol{Q}\in[0,BV+1]^{n}$.
2 for each iteration $t\in[1:T]$ do
3       Sample $a_{t}\sim\boldsymbol{p}(t)$ and set $\boldsymbol{Y}(t)=\boldsymbol{e}_{a_{t}}$ where $\boldsymbol{e}_{i}$ is the unit vector with $i$-th entry being 1.
4       Receive feedback $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$.
5       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using, $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[1:n]$.
6       Find, $\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1)$ by solving the problem,
$(\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1))=\arg\min_{\boldsymbol{p}\in\Delta_{n,\varepsilon},\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})$
$+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big{]},$ (29)
where $D_{\text{KL}}$ is the KL-divergence.
7       Update the virtual queues,
$\boldsymbol{Q}(t+1)=[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)]_{+},$ (30)
where $\boldsymbol{X}(t)=\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
8
Algorithm 2 Single Channel Adpative MAC
Most of the results required for the error analysis are borrowed from the multi-channel case. We only prove the results that are unique to this case. We first focus on solving the intermediate problem (6). Notice that the problem can be separated into the problem of finding optimal $\boldsymbol{\gamma}(t+1)$ and $\boldsymbol{p}(t+1)$.
Finding $\boldsymbol{\gamma}(t+1)$: This is the same problem as in the multi-channel case.
Finding $\boldsymbol{p}(t+1)$: This is (P3) with $\boldsymbol{x}=\eta\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)$, and $\boldsymbol{y}=\boldsymbol{p}(t)$.
Similar to Lemma 5, we have Pinsker’s inequality for KL divergence.
Lemma 12 (Pinsker’s inequality)
For $\boldsymbol{x},\boldsymbol{y}\in\Delta_{n}$, we have that,
$D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert_{1}^{2}\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert^{2},$
We also have the following lemma similar to Lemma 6.
Lemma 13
We have $D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\leq\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{x}\in\Delta_{n}$ where $\boldsymbol{y}\in\Delta_{n,\varepsilon}$.
Notice that since the queue updates and the problem to solve $\boldsymbol{\gamma}(t+1)$ are the same as in the multi-channel case, we have the same queue bound.
Lemma 14
We have that for all $t\in\{1,2,,\dots\}$ and $i\in\{1,2\dots,n\}$, $Q_{i}(t)\leq BV+1$.
Define the drift $\Delta(t)$ as,
$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}.$ (31)
We have the following lemma.
Lemma 15
We have that for all $t\in\{1,2,\dots\}$,
$\Delta(t)\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
Proof:
Proof follows repeating the same arguments as Lemma 8. ∎
Lemma 16
We have that,
-- $\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
$\ \ \ \ \leq\frac{\eta n}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$
where $\hat{S}_{i}(t)$ is defined in line 5 of Algorithm 2.
Proof:
Notice that,
$\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)[p_{i}(t+1)-p_{i}(t)]\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(a)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{2\eta}\mathbb{E}\left\{\lVert\boldsymbol{p}(t+1)-\boldsymbol{p}(t)\rVert^{2}\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(b)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$ (32)
where (a) follows from $\frac{1}{2\eta}\lVert\boldsymbol{a}\rVert^{2}+\frac{\eta}{2}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{n}a_{i}b_{i}$ for $n$-dimensional vectors $\boldsymbol{a,b}$, (b) follows from Lemma 12. Define the history $\mathcal{H}(t)$ similar to multi-channel case (see, (23)). Notice that $\boldsymbol{p}(\tau),\boldsymbol{Q}(\tau),\boldsymbol{\gamma}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable. Now, we handle the terms of (16) separately. Notice that,
$\mathbb{E}$$\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}^{2}(t)\right\}$
$=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}\hat{S}_{a_{t}}(t)\right\}\leq_{(a)}\frac{1}{\varepsilon}\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}(t)\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\right\}$
$=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=_{(b)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}\mathbb{E}\left\{S_{i}(t)Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}$
$=_{(c)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}q_{i}\mathbb{E}\left\{Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(d)}\frac{1}{\varepsilon}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
$\leq\frac{n}{\varepsilon}(BV+1)^{2}$
where $a_{t}$ is defined in line 3 of Algorithm 2, (a) follows since $S_{a_{t}}(t)\leq 1$ and $p_{a_{t}}(t)\geq\varepsilon$, (b) follows since $\boldsymbol{Q}(t)$ and $\boldsymbol{p}(t)$ are $\mathcal{H}(t)$-measurable, (c) follows since $\boldsymbol{S}(t)$ is independent of $\boldsymbol{Y}(t)$ and $\mathcal{H}(t)$, (d) follows since $q_{i}\leq 1$, and the last inequality follows due to Lemma 14. Also,
$\sum_{i=1}^{n}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}=\mathbb{E}\left\{Q_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}p_{a_{t}}(t)\right\}=\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
Combining everything and substituting in (16), we are done. ∎
Now, we introduce the following lemma.
Lemma 17
We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
$VT\phi(\boldsymbol{\gamma})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
Proof:
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 15, we have that,
$\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{\eta n}{2\varepsilon}(BV+1)^{2}$
$\ \ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$ (33)
where (a) follows by Lemma 16. Now notice that combining the decision (6) with Lemma 4, we have that for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))-\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1)).$
Taking expectations of the above, we have that,
$-V$ $\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ \ -\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}p_{i}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}.$
Substituting the above in (17), we have that,
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n+\frac{\eta n}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, we have that,
$\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\right\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right),$
where for the last inequality, we have used Lemma 13. Rearranging the above and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Now, we are ready to establish the performance bound of the algorithm.
Theorem 2
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
$\phi^{\text{opt}}$ $-\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
Proof:
Consider the problem
$\underset{\boldsymbol{p},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (34)
[…]
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (36)
$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (37)
Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is $\phi^{\text{opt}}$. Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote the solution of the above problem. Substituting $\boldsymbol{p}^{*}(1-\varepsilon n)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 17 we have that,
$VT\phi^{\text{opt}}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i}(1-\varepsilon n)-\varepsilon q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}nq_{i}p^{*}_{i}\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
$\leq_{(a)}nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\varepsilon n^{2}T(BV+1)+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (38)
where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (34)-(37)), and Lemma 18. Now, we use the Jensen’s inequality to obtain,
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
Notice that Lemma 11 holds since the queue updates are the same as in the multi-channel case. Combining above with Lemma 11, we have that,
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Lemma 18. Now, using Jensen’s inequality, we are done. ∎
Comparing the error bounds of Theorem 2 and Theorem 1 (with $m=1$), it can be seen that the bound of Theorem 2 is strictly better. In particular, the second term of the error bound has a $\Theta(n)$ dependence on $n$ in Theorem 2, whereas the corresponding value is $\Theta(n^{3})$ in Theorem 1.
V UCB Algorithm
In this section, we present a UCB-based algorithm for the link selection problem in MAC (Algorithm 3), after which we move on to the analysis. Recall that $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ is the user-channel assignment during time slot $t$. For $t\in[s]$, we set $\boldsymbol{Y}(t)$ arbitrarily such that each user-channel pair is explored exactly once during the interval $[s]$. For each $t>s$, we compute a permutation matrix $\boldsymbol{M}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1. The assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ for the $t>s$ is obtained by discarding the last $s-m$ columns or the last $s-n$ rows of $\boldsymbol{M}(t)$ depending on whether $n$ or $m$ is larger. The algorithm uses parameters $\delta_{s},\delta_{s+1},\dots$ that satisfy $\delta_{t}>0$ for all $t\in\{s,s+1,\dots\}$. For each $i\in[n]$, $j\in[m]$, we define,
$n_{i,j}(t)=\sum_{\tau=1}^{t}Y_{i,j}(t),\ \forall t\in\mathbb{N}$
$\hat{S}_{i,j}(t)=\begin{cases}\frac{\sum_{\tau=1}^{t}Y_{i,j}(t)S_{i,j}(t)}{n_{i,j}(t)}&\text{ if }n_{i,j}(t)>0\\ 0&\text{ otherwise}\end{cases},\ \forall t\in\mathbb{N}$
$f_{i,j}(t)=\sqrt{\frac{\log\left(\frac{n_{i,j}(t)[n_{i,j}(t)+1]}{\delta_{t}}\right)}{2n_{i,j}(t)}},\ \forall t\in\{s,s+1,\dots\}$
$\text{UCB}_{i,j}(t)=\hat{S}_{i,j}(t)+f_{i,j}(t),\ \forall t\in\{s,s+1,\dots\}$ (39)
1 For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
2
3Initialize $\boldsymbol{Q}(s+1)=0$.
4 for each time slot $t\in\{s+1,s+2,\dots,\}$ do
5       Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{M}(t)$ by solving
$\boldsymbol{\gamma}(t)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{N}Q_{i}(t)\gamma_{i},$ (40)
$\boldsymbol{M}(t)=\arg\max_{\boldsymbol{M}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
where $\text{UCB}_{i,j}(t-1)$ is defined in (V).
6       Use the assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$, where $Y_{i,j}(t)=M_{i,j}(t)$ for $(i,j)\in[n]\times[m]$ and receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
7       Update the queues
$Q_{i}(t+1)=[Q_{i}(t)+\gamma_{i}(t)-X_{i}(t)]_{+},$ (41)
$\forall\ i\in[n]$, where $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$.
8
Algorithm 3 UCB MAC
We first focus on solving the inner problem (40). Notice that the problem to solve to obtain $\boldsymbol{\gamma}(t)$ is the same as for Algorithm 1. The problem to solve to obtain $\boldsymbol{M}(t)$ has a classic max-weight structure (hence, $\boldsymbol{M}(t)$ is a permutation matrix). Hence, we can use the Hungarian algorithm [49] to obtain the solution.
Now, we move on to the analysis of the Algorithm 3. Since the queue updates and the problem to solve to obtain $\boldsymbol{\gamma}(t)$ are the same as in Algorithm 1, we have the same deterministic queue bound. We formally state this in the following lemma.
Lemma 18
We have for all $t\in\{s+1,s+2,\dots\}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$
Proof:
This follows repeating the argument of Lemma 7. ∎
Define the collection of good events $\Omega_{s+1},\Omega_{s+2},\dots$ as
[…]
where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (V). We have the following lemma.
Lemma 19
For each $t\in\{s+1,s+2,\dots,\}$, we have $\mathbb{P}\{\Omega_{t}\}\geq 1-2nm\delta_{t-1}$.
Proof:
The proof follows directly by applying the Hoeffding inequality with union bound. See [48] for more details. ∎
Define the drift $\Delta_{t}$
$\Delta_{t}=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}-\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
and history $\mathcal{H}(t)$ as
[…]
for each $t\in\{s+1,s+2,\dots\}$. Notice that $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable.
Lemma 20
We have for each $t\in\{s+1,s+2,\dots\}$
$\Delta(t)\leq n+\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\right\}.$
Proof:
Proof is similar to the proof of Lemma 8 and is given in Appendix B. ∎
Fix a time horizon $T\in\{s+1,s+2,\dots\}$. Then we have the following result.
Lemma 21
Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{\gamma}(t)\right)\right\}$ $\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}$
$\ \ \ \ +\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t},$
where $f_{i,j}(t)$ is defined in (V).
Proof:
We begin with the following two claims.
Claim 1: We have that,
$-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\leq V\phi^{\text{max}}+nBV+n$
Proof:
The proof immediately follows from Lemma 18 and the definition of $\phi^{\text{max}}$ in Assumption A1. ∎
Claim 2: We have that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\Bigg{|}\Omega_{t}\right\}$
$\leq-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
Proof:
From the definition of the good event $\Omega_{t}$ in (42), we have
$\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\geq\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}-2\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\geq\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)M_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\ \ \ \ -2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$ (44)
where the last inequality follows from Lemma 18. Also, notice that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}(t)\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)M_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\leq_{(a)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}$
$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}\leq-V\phi^{\text{opt}}$ (45)
where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{M}(t)$ for intermediate problem (40) in Algorithm 3 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (42); the last inequality follows from (11). Adding the inequalities (21) and (21) and rearranging, we are done. ∎
Combining claim 1 and claim 2, and the law of total probability, we have that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\right\}$
$\leq(V\phi^{\text{max}}+nBV+n)\mathbb{P}\{\Omega_{t}^{c}\}-V\phi^{\text{opt}}\mathbb{P}\{\Omega_{t}\}$
$\ \ \ \ +2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}\mathbb{P}\{\Omega_{t}\}$
$\leq_{(a)}(V\phi^{\text{max}}+V\phi^{\text{opt}}+nBV+n)\mathbb{P}\{\Omega_{t}^{c}\}-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\leq_{(b)}2(2V\phi^{\text{max}}+nBV+n)nm\delta_{t-1}-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$ (46)
where (a) follows from the fact that for a nonnegative random variable $\boldsymbol{X}$ and an event $\mathcal{H}$ we have $\mathbb{E}\{X|\mathcal{H}\}\mathbb{P}\{\mathcal{H}\}\leq\mathbb{E}\{X\}$ and (b) follows since $\phi^{\text{opt}}\leq\phi^{\text{max}}$ and Lemma 19. Adding (21) to the result of Lemma 20 and rearranging, we have
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}+2(2V\phi^{\text{max}}+nBV+n)nm\delta_{t-1}$
Now, we sum the above for $t\in[s+1,T]$ to obtain
$\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T+1)\rVert^{2}-\lVert\boldsymbol{Q}(s+1)\rVert^{2}\}-V\sum_{t=s+1}^{T}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n(T-s)-V(T-s)\phi^{\text{opt}}+2(BV+1)\sum_{t=s+1}^{T}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\ \ \ \ +2(2V\phi^{\text{max}}+nBV+n)nm\sum_{t=s}^{T-1}\delta_{t}.$
Using $\lVert\boldsymbol{Q}(T+1)\rVert^{2}\geq 0$ and $\lVert\boldsymbol{Q}(s+1)\rVert^{2}=0$, we have
$V(T-s)\phi^{\text{opt}}-V\sum_{t=s+1}^{T}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n(T-s)+2(BV+1)\sum_{t=s+1}^{T}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\ \ \ \ +2(2V\phi^{\text{max}}+nBV+n)nm\sum_{t=s}^{T-1}\delta_{t}.$
Dividing both sides by $V(T-s)$, and using Jensen’s inequality, we are done. ∎
Now we have the following lemma that combines $\boldsymbol{\gamma}(s+1),\boldsymbol{\gamma}(s+2),\dots,\boldsymbol{\gamma}(T)$ with $\boldsymbol{X}(s+1),\boldsymbol{X}(s+2),\dots,\boldsymbol{X}(T)$.
Lemma 22
We have
$\phi\left(\frac{1}{T-s}\left(\sum_{t=s+1}^{T}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T-s}\left(\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)\right)+\frac{nB(BV+1)}{T-s}.$
Proof:
The proof uses the same argument as Lemma 11. ∎
Now, we have the following theorem that establishes the performance bound of Algorithm 3.
Theorem 3
Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
$\phi^{\text{opt}}$ $-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (47)
where $f_{i,j}(t)$ is defined in (V). In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
  1. 1.
    For all $V>0$,
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
    $\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{V(T-s)}$
    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
  2. 2.
    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right),$
    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
  3. 3.
    Fix $\epsilon>0$. Using $V=\Theta(1/\epsilon)$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}(\epsilon),$
    and $\mathcal{O}$ hides the dependence on all parameters but $\epsilon$.
Proof:
First, notice that
$\left\lvert\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)-\phi\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\rvert$
$\leq\sqrt{n}B\left\lVert\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)-\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\rVert\sqrt{n}B\left\lVert\frac{1}{T}\sum_{t=1}^{s}\boldsymbol{X}(t)+\frac{s}{T(T-s)}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right\rVert$
$\leq\frac{\sqrt{n}B}{T}\sum_{t=1}^{s}\left\lVert\boldsymbol{X}(t)\right\rVert+\frac{s}{T(T-s)}\sum_{t=s+1}^{T}\left\lVert\boldsymbol{X}(t)\right\rVert\leq\frac{\sqrt{n}sB}{T}+\frac{s}{T}=\frac{(\sqrt{n}B+1)s}{T},$ (48)
where the first inequality follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.
Combining Lemma 22 and Lemma 21, we have
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)\right\}$
$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}.$ (49)
Combining (3) and (3) and using the Jensen’s inequality, we have (3).
To prove part 1, first fix $i\in[n]$ and $j\in[m]$. When $\delta_{t-1}=1/(t-1)$ for all $t\in[s+1:T]$, we have
$f_{i,j}(t-1)$ $=\sqrt{\frac{\log\left(\frac{n_{i,j}(t-1)[n_{i,j}(t-1)+1]}{\delta_{t-1}}\right)}{2n_{i,j}(t-1)}}\leq\sqrt{\frac{3\log\left(T\right)}{2n_{i,j}(t-1)}}.$ (50)
Also, $n_{i,j}(t)-n_{i,j}(t-1)=1$ if and only if $Y_{i,j}(t)=1$. Hence, we have
$\mathbb{E}\left\{\sum_{t=s+1}^{T}\frac{Y_{i,j}(t)}{\sqrt{n_{i,j}(t-1)}}\right\}=\mathbb{E}\left\{\sum_{k=1}^{n_{i,j}(T-1)-1}\sqrt{\frac{1}{k}}\right\}\leq\mathbb{E}\left\{\sum_{k=1}^{T}\sqrt{\frac{1}{k}}\right\}\leq 2\sqrt{T}$ (51)
where for the last inequality we have used $\sum_{\tau=1}^{t}\frac{1}{\sqrt{\tau}}\leq 2\sqrt{t}$ for all $t\geq 1$. Combining (50) and (51), we have
$\sum_{t=s+1}^{T}\mathbb{E}\{Y_{i,j}(t)f_{i,j}(t-1)\}\leq\sqrt{6T\log(T)}$
Now substituting the above in (3), we have
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\frac{1}{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$ (52)
Part 1 follows from above after using the fact that $\sum_{\tau=1}^{t}\frac{1}{\tau}\leq\ln(t)+1$ for all $t\geq 1$.
For part 2, notice that since $T\geq 2s$, from part 1 we have
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{4\sqrt{6}nm(BV+1)\sqrt{\log(T)}}{V\sqrt{T}}+\frac{4(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{VT}$
$\ \ \ \ +\frac{2nB(BV+1)}{T}+\frac{(\sqrt{n}B+1)s}{T},$
which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$.
For part 3, notice that from part 1,
$\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)\leq\frac{n}{V}.$
Using $V=\Theta(1/\epsilon)$, we are done. ∎
VI Discussion
VI-A Distributed Implementation
Consider a system where the $n$ users take decisions in the absence of a centralized controller. In each time slot, each user receives as feedback whether each channel was successfully accessed or not (this can be, for instance, via sensing each channel). Notice that $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$ are $\mathcal{H}(t)$-measurable, where $\mathcal{H}(t)$ is defined in (43). Hence, during time slot $t$, if each user has access to the full history $\mathcal{H}(t)$, they can independently implement Algorithm 3 to find $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$. From the feedback, each user can compute $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$ and hence the full history $\mathcal{H}(t+1)$. Hence, starting from $\mathcal{H}(s+1)$, each user can independently implement Algorithm 3. Notice that this is not possible in Algorithm 1, since the channel assignments are not $\mathcal{H}(t)$-measurable.
VII Simulations
For simulations, we consider four scenarios with $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+x_{i})$, $n=5,m=3$, and $T=10^{5}$. In each scenario, the values of $q_{i,j}$ change mid-way through the simulation (at $T=5\times 10^{4}$). In Scenario 1 (Figure 1-Left-Top), certain user-channel links are turned off at $T=5\times 10^{4}$. In Scenario 2 (Figure 1-Right-Top), certain user-channel links that were turned off before, turn on at $T=5\times 10^{4}$. In Scenarios 3 (Figure 1-Left-Bottom) and 4 (Figure 1-Right-Bottom), the values of $q_{i,j}$ for certain user-channel links change at $T=5\times 10^{4}$. In each scenario, the algorithms are not told about the change. For each figure, for each algorithm, we plot $\phi\left(\frac{1}{t}\sum_{\tau=1}^{t}\boldsymbol{X}(\tau)\right)$ for $t\leq T/2$ and $\phi\left(\frac{1}{t-T/2}\sum_{\tau=T/2+1}^{t}\boldsymbol{X}(\tau)\right)$ for $t>T/2$ vs $t$. We also plot the optimal objective value of (P2), for reference. Note that in the first half of the simulation, both algorithms converge to the optimal solution in all scenarios, although the convergence of Algorithm 3 is faster. However, to show the importance of Algorithm 1, we observe the second half of the simulation. Although Algorithm 3 achieves superior performance in the first half of the simulation, it does not perform well in the second half except in Scenario 2. In contrast Algorithm 1 achieves similar performance in both halves in all scenarios. This is due to the adaptiveness of Algorithm 1.
Figure 1: Top-Left: Scenario 1. Top-Right: Scenario 2. Bottom-Left: Scenario 3. Bottom-Right: Scenario 4.
VIII Conclusions
This paper focused on the problem of designing algorithms for automatic link selection in multi-channel multiple access with link failures. In particular, we solved a network utility maximization problem with matching constraints and bandit feedback on link failures. We considered two algorithms, where the first algorithm has slower convergence and is adaptive, and the second algorithm has faster convergence and is not adaptive.
Appendix A Proof of Lemma 2
Before proving this, notice that due to the normalization steps 1 and 2 of the ROUND function, we have $\boldsymbol{1}\geq\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}$, and $\boldsymbol{1}\geq(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}$, where the inequalities are taken entrywise. Hence, we have $\boldsymbol{Q}\geq\boldsymbol{P}^{{}^{\prime\prime}}\geq 0$. Now, we prove each part separately.
1) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then $\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{row}}$ and $\boldsymbol{Q}=\boldsymbol{P}^{{}^{\prime\prime}}$ (by the definition of $\boldsymbol{Q}$ in (19)). Also, due to the normalization of the columns in step 2, each column of $\boldsymbol{P}^{{}^{\prime\prime}}$ has a sum of at most 1. However, notice that since $\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{row}}$, the sum of its entries is $n$. Hence, the sum of each column must be exactly 1. Hence, $\boldsymbol{Q}=\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{doub}}$ as desired. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Then we have
$\boldsymbol{Q}\boldsymbol{1}$ $=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\boldsymbol{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=_{(a)}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}})\boldsymbol{1}}{\boldsymbol{1}^{\top}(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{s\boldsymbol{1}-s\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(s-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})}{s-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}=\boldsymbol{1},$
where in (a) we have used $\boldsymbol{1}\geq\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}$ when simplifying the denominator. The claim $\boldsymbol{1}^{\top}\boldsymbol{Q}=\boldsymbol{1}^{\top}$ follows repeating the same argument.
2) We will prove the case $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$. The other case follows similarly. Since $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$, from (17), we have $P^{{}^{\prime}}_{i,j}=P_{i,j}$. Also, since $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$ we have $P_{i,j}\leq 1$ for all $i,j\in\{1,\dots,s\}$. Hence $P^{{}^{\prime}}_{i,j}\leq 1$ for all $i,j\in\{1,\dots,s\}$. Hence, from (18), we have $P^{{}^{\prime\prime}}_{i,j}\geq P{{}^{\prime}}_{i,j}/s=P_{i,j}/s\geq\varepsilon/s$. Since we already know $\boldsymbol{Q}\geq\boldsymbol{P}^{{}^{\prime\prime}}$, we have $Q_{i,j}\geq\varepsilon/s$ for all $i,j$. Part 1 shows $Q\in\mathcal{S}^{\text{doub}}$, so $Q\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$. So we are done.
3) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0$. Also, due to the scaling, we have $s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\geq 0$. Hence, we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ and we are done. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Notice that
$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=_{(a)}\frac{\boldsymbol{1}^{\top}\left(\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\right)\boldsymbol{1}}{\boldsymbol{1}^{\top}\left(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\right)}$
$=\frac{\boldsymbol{1}^{\top}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\frac{s^{2}-2s\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}^{2}}{s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}}=s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1},$
where (a) follows since $\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\geq 0$, and $\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}\geq 0$, as a result of which $(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\geq 0$.
4) Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s$
$=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
$=\sum_{i\in\mathcal{U}}\left(\sum_{j=1}^{s}P_{i,j}-1\right)+\sum_{i\in\mathcal{U}^{c}}\left(1-\sum_{j=1}^{s}P_{i,j}\right)+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
$=2\sum_{i\in\mathcal{U}}\sum_{j=1}^{s}P_{i,j}-|\mathcal{U}|+|\mathcal{U}^{c}|-s$
$=2\left(\sum_{i\in\mathcal{U}}\sum_{j=1}^{s}P_{i,j}-|\mathcal{U}|\right)=2\sum_{i\in\mathcal{U}}\left(\sum_{j=1}^{s}P_{i,j}-1\right)$
$=_{(a)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime}}\rVert_{1})=2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2(\lVert\boldsymbol{P}^{{}^{\prime}}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})$
$=_{(b)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left[\sum_{i=1}^{s}P^{{}^{\prime}}_{i,j}-1\right]_{+}\right)$
$\geq_{(c)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left[\sum_{i=1}^{s}P_{i,j}-1\right]_{+}\right)$
$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}P_{i,j}-1\right\rvert\right)$
$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$ (53)
where (a) and (b) follow from the definitions of $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ in (17) and (18), respectively; (c) follows since $P_{i,j}\geq P^{{}^{\prime}}_{i,j}$ due to scaling. Rearranging the above inequality, we have
$\lVert\boldsymbol{P}\rVert_{1}-2\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+s\leq\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}.$ (54)
Now, to complete the proof notice that
$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
$=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
$\leq_{(b)}\lVert\boldsymbol{P}\rVert_{1}-2\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+s\leq_{(c)}\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$
$\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right),$ (55)
where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to part 3, and (c) follows from (54).
Appendix B Proof of Lemma 20
Notice that from the queuing equation (41), we have
$Q^{2}_{i}(t+1)$ $\leq(Q_{i}(t)+\gamma_{i}(t)-X_{i}(t))^{2}\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t)+X^{2}_{i}(t)+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
for all $i\in[n]$. Summing the above for $i\in[n]$, we have
$\lVert\boldsymbol{Q}(t+1)\rVert^{2}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)].$
Taking the expectations conditioned on the history $\mathcal{H}(t)$ defined in (43), we have
$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\mathbb{E}\{\gamma_{i}(t)|\mathcal{H}(t)\}-\mathbb{E}\{X_{i}(t)|\mathcal{H}(t)\}]$
$=\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right],$
where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result.

### v2

@@
 This paper focuses on the problem of automatic link selection in multi-channel multiple access control using bandit feedback.
-In particular, a controller assigns multiple users to multiple channels in a time slotted system, where in each time slot at most one user can be assigned to a given channel and at most one channel can be assigned to a given user.
-Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed probability $f_{i,j}$.
-The failure probabilities are not known to the controller.
+In particular, a controller assigns multiple users to multiple channels in a time-slotted system, where in each time slot, at most one user can be assigned to a given channel, and at most one channel can be assigned to a given user.
+Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed unknown probability $1-q_{i,j}$.
 The assignments are made dynamically using success/failure feedback.
-The goal is to maximize the time average utility, where we consider an arbitrary (possibly nonsmooth) concave and entrywise nondecreasing utility function.
-The problem of merely maximizing the total throughput has a solution of always assigning the same user-channel pairs and can be unfair to certain users, particularly when the number of channels is less than the number of users.
-Instead, our scheme allows various types of fairness, such as proportional fairness, maximizing the minimum, or combinations of these by defining the appropriate utility function.
-We propose two algorithms for this task.
-The first algorithm is adaptive and gets within $\mathcal{O}(\log(T)/T^{1/3})$ of optimality over any interval of $T$ consecutive slots over which the success probabilities do not change.
-The second algorithm has faster $\mathcal{O}(\sqrt{\log(T)/T})$ performance over the first $T$ slots, but does not adapt well if probabilities change.
-Index Terms:
-Multi-armed bandit learning; Proportional fairness; Network utility maximization; Optimization; Stochastic control; Adaptive learning
+The goal is to maximize the time-average utility, where we consider an arbitrary (possibly nonsmooth) concave, entrywise nondecreasing utility function.
+The first proposed algorithm has fast $\mathcal{O}(\sqrt{\log(T)/T})$ convergence.
+However, this algorithm requires solving a convex optimization problem within each iteration, which can be computationally expensive.
+The second algorithm has slower $\mathcal{O}(\sqrt[3]{\log(T)/T})$ convergence, while avoiding the costly inner optimization.
+Both of these algorithms are adaptive.
+In particular, the convergence guarantee holds for any interval of $T$ consecutive slots during which the success probabilities do not change.
+We further study several special cases.
+In the single-channel setting, we obtain both fast $\mathcal{O}(\sqrt{\log(T)/T})$ convergence and efficient implementation via a simpler adaptive mechanism.
+We also consider a UCB-based non-adaptive algorithm with max-weight-type decisions.
+Simulations highlight intriguing performance trade-offs and demonstrate rapid adaptation of the proposed adaptive schemes.
+Automatic Link Selection in Multi-Channel Multiple Access with Link Failures
+Mevan Wijewardena, Michael J.
+Neely, Haipeng Luo Mevan Wijewardena and Michael J.
+Neely are with the Electrical Engineering department at the University of Southern California.
+Haipeng Luo is with the Computer Science department at the University of Southern California.
+This work was presented in part at the WiOpt conference in Linköping, Sweden, May 2025 [50].
 I Introduction
 We consider the Multiple Access Control (MAC) problem with $n$ users and $m$ channels in slotted time $t\in\mathbb{N}$.
-In each time slot, a controller has to assign the users to channels such that at most one user is assigned to given channel and at most one channel is assigned to a given user.
-The channel assignments may fail.
-In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the assignment fails independently with probability $1-q_{i,j}$.
+In each time slot, a controller assigns users to channels such that at most one user is assigned to a given channel and at most one channel is assigned to a given user.
+The transmissions over each channel may fail.
+In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the transmission fails independently with probability $1-q_{i,j}$.
 The controller does not know the probabilities $q_{i,j}$.
-Instead, at the end of every slot, it receives feedback on whether the transmission for each assigned user-channel pair succeeded or failed.
+Instead, at the end of every slot, it receives feedback on the successes.
 Define the matrices $\boldsymbol{Y}(t),\boldsymbol{S}(t)\in\{0,1\}^{n\times m}$ and vector $\boldsymbol{X}(t)\in\{0,1\}^{n}$, where
-$S_{i,j}(t)=\begin{cases}1&\text{ if link }i,j\text{ is successful in time slot }t\\ 0&\text{ otherwise},\end{cases}$
-$Y_{i,j}(t)=\begin{cases}1&\text{ user }i\text{ is assigned to channel }j\text{ in time slot }t\\ 0&\text{ otherwise},\end{cases}$
+[…]
+[…]
 and $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for all $i\in\{1,2,\dots,n\}$.
-The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the link failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.
+The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the transmission failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.
 [^1] The limit is assumed to exist for simplicity of this introduction; the precise goal is to maximize a $\lim\inf_{T\to\infty}\phi(\boldsymbol{\overline{X}}(T))$
@@
 $\lim_{T\to\infty}g(T)=0$
-In addition, we are looking for algorithms that are adaptive. Formally, consider a system in which the channel success probabilities may change. In such a system, given a $T\in\mathbb{N}$, we require
-$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$
-for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $[T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame.
-Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $[T_{0}:T_{0}+T-1]$ of the above scenario.
+In addition, we are looking for algorithms that are adaptive.
+Formally, consider a system in which the channel success probabilities $q_{i,j}$ may change.
+In such a system, given a $T\in\mathbb{N}$, we require
+$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$ (1)
+for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame.
+Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $T_{0}:T_{0}+T-1]$ of the above scenario.
 Note that $g$ is the same function regardless of $T_{0}$.
+In fact, it can be shown that given $T^{{}^{\prime}}\in\mathbb{N}$, our algorithms can achieve
+$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T^{{}^{\prime}})$
+for any $T_{0}\in\mathbb{N}$ and $T\geq T^{{}^{\prime}}$, where the algorithm parameters only depend on $T^{{}^{\prime}}$ (not on $T$ or $T_{0}$).
+Hence, given an error tolerance $\epsilon>0$, we can find $T^{{}^{\prime}}$ such that $g(T)\leq\epsilon$ for all $T\geq T^{{}^{\prime}}$.
+This will ensure adaptation in the case where channel probabilities remain fixed for windows longer than $T^{{}^{\prime}}$.
 I-A Utility functions
@@
 However, if there is one user with very low success probability, this utility function can cause almost all the resources to be devoted to that user, resulting in poor performance for all users.
-Another choice is $\phi(\boldsymbol{x})=w_{1}\min\{x_{1},\dots,x_{n}\}+w_{2}\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $w_{1},w_{2},\beta$ are given nonnegative weights.
+Another choice is $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta$ are given nonnegative weights.
 The logarithmic term introduces a form of proportional fairness [29, 30].
@@
 Some examples include scheduling in wireless networks [23, 28], computer networks [37], vehicular networks [14] and unmanned aerial vehicle networks [31].
-These works consider different goals such as optimal scheduling for queue-stability [11], power minimization [12], utility maximization [13], ensuring fairness [14], flow control [15], and multiple access control [16].
+These works consider different goals such as optimal scheduling for queue-stability [27], power minimization [18], utility maximization [43].
 Multiple access, where many users access a limited number of communication channels, is an important problem in network scheduling.
 Here, it is desirable for users to be scheduled to avoid collisions.
-Link failures occur when the receiver is unable to decode the packet transmissions.
+Transmission failures occur when the receiver is unable to decode the packet transmissions.
 This can occur, for example, when a fixed transmission power is used, but channel conditions have random and unknown fluctuations so that the received signal strength is insufficient for decoding.
 Different links can have different properties (such as different geographic distances to the receiver), so they can also have different success probabilities.
-The naive approach of assigning the users with links with the least failure probabilities leads to unfairness since, in such a scenario, users with high link failure probabilities will never be assigned.
+The naive approach of assigning the users to links with the least failure probabilities leads to unfairness since, in such a scenario, users with high transmission failure probabilities will never be assigned.
 A common approach to solve this problem is to maximize a utility function of the time-averaged success.
@@
 The case when success probabilities are known in advance can be solved offline as a convex optimization problem using the mirror descent technique [20].
-The problem becomes challenging when the success probabilities are not known, but we only receive bandit feedback on the successes.
-The problem has to be approached by combining ideas from optimizing functions of time averages with multi-armed bandit learning.
-The work on bandits with vector rewards and concave utility functions can be adapted for the single channel case ($m=1$) of our problem (See [26, 27, 28]).
-There are two main drawbacks to these approaches.
-First, the above works do not consider the matching constraints considered in our work.
-Next, they focus on upper confidence bound (UCB) techniques and are not adaptive.
-It is possible to develop adaptive algorithms using techniques from the above works together with the EXP3 algorithm [29] for the single channel case.
-However, they cannot be directly extended for the multi-channel case.
-The main reason for this is the complexity of the inner problem arising in each iteration which is a problem over the set of doubly stochastic matrices (Birkhoff polytope).
-This can be addressed using Sinkhorn’s algorithm [30] in each iteration.
-However, this results in an algorithm that requires computationally complex inner iterations to be performed within each iteration.
-The work of [31] uses follow the regularized leader approach to solve adversarial bandit problems over the set of doubly stochastic matrices.
-However, their algorithm relies on a computationally expensive inner iteration similar to what we described before.
-The work of  [32] proposes an algorithm to solve online optimization problems over transport polytopes.
-All the inner iterations of their algorithm have explicit solutions thanks to the rounding trick introduced for transport polytopes by [33].
-In our work, we adapt the rounding trick to the Birkhoff polytope and develop an algorithm for general $m$.
-We also develop a UCB-based algorithm for our problem.
-It should be noted that although this algorithm achieves faster convergence, it is not adaptive.
-However, the UCB algorithm has additional advantages such as distributed implementation.
-Our paper also treats general (possibly nonsmooth) utility functions, so Frank-Wolfe methods that rely on smoothness cannot be used.
-Existing work on using learning for scheduling problems in communications include spectrum sharing [34, 35], network utility maximization with unknown utility functions [36], and queue stability [11].
-The existing literature on bandit learning-based methods for scheduling problems focuses on maximizing throughput [37].
-The work of [36] considers bandit learning with unknown utilities, and the work of [38] considers the online resource allocation problem, where the objective is to maximize the $\alpha$-fairness of the allocations.
-Neither of the above works considers the bandit reward structure, and both works have a different problem structure from ours.
-In addition, the above works do not focus on the adaptiveness of the proposed algorithms.
+The problem becomes challenging when the success probabilities are unknown and only bandit feedback on transmission successes is available.
+As a result, the problem must be addressed by combining ideas from utility optimization over time averages with multi-armed bandit learning.
+Existing work on bandits with vector rewards and concave utility functions can be adapted to the single-channel case ($m=1$) of our problem (see [2, 19, 3]).
+However, these approaches have two main drawbacks.
+First, they do not consider the matching constraints present in our setting.
+Second, they rely on upper confidence bound (UCB) techniques and are not adaptive.
+In particular, these algorithms cannot adapt to changes in transmission failure probabilities without being explicitly informed of such changes.
+Hence, the goal of this paper is to develop adaptive algorithms for utility optimization with matching constraints under bandit feedback.
+This is achieved by combining the EXP3.S algorithm [7], which enables adaptive bandit learning, with Lyapunov optimization [42] to address the utility maximization objective.
+Although EXP3.S was originally developed for adversarial bandits, it has also been successfully applied in nonstationary stochastic environments [27].
+Existing work on using learning for scheduling problems in communications includes spectrum sharing [11, 21], network utility maximization with unknown utility functions under full feedback [25], and queue stability [27].
+The literature on bandit learning-based scheduling primarily focuses on maximizing throughput [46].
+The work of [25] considers network utility maximization where the utility functions are learned online, and [44] studies online resource allocation with the objective of maximizing $\alpha$-fairness.
+Neither of these works considers a bandit reward structure or matching-type constraints, and both differ from our problem formulation.
+The work of [10] studies an online optimization problem on transport polytopes.
+While their framework can be adapted to the Birkhoff polytope to incorporate matching constraints, it does not extend to bandit feedback settings.
+Moreover, none of the above works explicitly addresses the adaptiveness of the proposed algorithms.
 It should be noted that the concept of adaptiveness considered in this paper is different from the adversarial setting [8, 48, 7], although they have similarities.
-In adversarial settings, optimality is defined with respect to the rewards received throughout the time horizon.
+In adversarial settings, the guarantee of convergence is defined with respect to the rewards received throughout the time horizon.
 This is useful when no stochastic assumptions can be placed on the rewards.
@@
 Here, the regret is modified to account for the changing environments and the regret bounds are in terms of some measure that captures the degree of change.
-Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [41, 42], and sliding window-based algorithms [43].
+Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [9, 16], and sliding window-based algorithms [26, 17].
 We utilize a simpler notion of adaptiveness and develop an algorithm for the case with general utility functions.
-Multi-armed bandit (MAB) learning [44, 45] is a class of extensively studied problems.
-In the classical MAB setting, a user chooses from multiple arms, each incurring a fixed mean cost.
-The choice is made without knowledge of the mean costs, using the feedback received on the cost of the chosen arm.
-The goal is to learn the arm with the lowest mean cost as fast as possible.
-Typically, the algorithms implemented to solve such problems involve an exploration phase, where the user explores all arms in order to learn the mean costs, after which, in the exploitation phase, the user exploits the learned information to make good decisions [46, 47].
-Upper confidence bound-based algorithms, where the algorithm maintains an upper bound on the mean cost of each arm, is a popular approach to solve these problems [47, 48].
-Adversarial bandit learning [39, 40], where an adversary is allowed to assign the cost of each arm after the user selects the arm, is an extension of multi-armed bandit learning.
-In such problems, no stochastic assumptions are placed on the rewards.
-These problems are solved using algorithms such as the EXP3 algorithm [29], where in each iteration of the algorithm, a probability distribution over the space of all possible actions is found, after which the decision is randomly sampled from the distribution.
-There are also many extensions of the MAB problem, such as linear bandits, contextual bandits, and combinatorial bandits [47].
+Multi-armed bandit (MAB) learning [34, 6] provides a principled framework for sequential decision-making under uncertainty.
+Classical stochastic bandit models focus on learning unknown mean rewards, often using upper confidence bound (UCB)–type algorithms [13, lattimore_szepesvári_2020].
+In contrast, adversarial bandit models make no stochastic assumptions on rewards and are commonly addressed using randomized algorithms such as EXP3 [7].
+Beyond these settings, numerous extensions have been studied, including linear, contextual, and combinatorial bandits [13].
+Table I positions our work relative to the most closely related literature.
+To the best of our knowledge, no prior work simultaneously addresses utility optimization under bandit feedback with matching constraints while providing interval-based adaptiveness guarantees.
+Unlike classical bandit algorithms, which maximize scalar time-averaged reward over independent arms, our setting involves combinatorial matching constraints and optimization of a concave utility function of the time-average throughput vector.
+These structural features introduce nonlinear coupling across users and channels and prevent a direct application of standard bandit techniques.
+Furthermore, achieving adaptiveness under changing channel statistics requires incorporating EXP3.S-style exploration mechanisms within a Lyapunov optimization framework.
+TABLE I: Comparison of work on Network Utility Optimization and Adaptive Bandits.
+Work Utility Feedback Environment Matching Adaptive Bound
+Constraints
+Adaptive MAC General Lipschitz Bandit Stochastic Yes Yes $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
+(This Paper)
+Adaptive
+MAC.CF General Lipschitz Bandit Stochastic Yes Yes $\mathcal{O}\left(\sqrt[3]{\frac{\log(T)}{T}}\right)$
+(This Paper)
+UCB MAC General Lipschitz Bandit Stochastic Yes No $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
+(This Paper)
+Neely [40] General Lipschitz Full Stochastic No Yes $\mathcal{O}\left(\sqrt{\frac{1}{T}}\right)$
+Agarwal et al [2] General Lipschitz Bandit Stochastic No No $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
+Neely [41] General Smooth Full Stochastic No No $\mathcal{O}\left(\frac{\log(T)}{T}\right)$
+Garivier et al [26] Linear Bandit Stochastic No Yes $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$
+Ballu et al [10] General Lipschitz Full Stochastic Yes No $\mathcal{O}\left(\sqrt{\frac{1}{T}}\right)$
 I-C Our Contributions
-We develop and analyze two algorithms to solve the problem of automatic link selection in multi-channel multiple access, combining the ideas of multi-armed bandit learning and Lyapunov optimization.
-Although the classical MAB problem has been widely analyzed for maximizing linear utilities, they suffer from lack of fairness in assignments when applied to the considered problem.
-It is notable that our method allows either smooth or nonsmooth concave, entrywise nondecreasing utilities.
-We prove that the first algorithm gets within $\mathcal{O}(T^{-1/3}\log(T))$ of optimality over any interval of $T$ consecutive time slots during which the (unknown) success probabilities do not change.
-If these probabilities are different before $T_{0}$, but change to new probabilities during $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, our performance guarantees for the new interval are independent of behaviors before $T_{0}$, even though the algorithm does not know the exact time $T_{0}$ of the change.
+  1. 1.
+    Novel Problem Formulation: We develop and analyze algorithms to solve the problem of automatic link selection in multi-channel multiple access by combining ideas from multi-armed bandit learning and Lyapunov optimization.
+Although the classical multi-armed bandit (MAB) problem has been widely studied for maximizing linear utilities, such approaches can suffer from a lack of fairness in assignments when applied to the problem considered here.
+In contrast, our framework allows for smooth or nonsmooth concave utility functions.
+  2. 2.
+    Adaptive Algorithm with Performance Guarantees: We develop two adaptive algorithms that jointly address bandit learning, matching constraints, and utility maximization.
+The algorithms combine the EXP3.S algorithm for adversarial bandits [7] with Lyapunov optimization and techniques for online matching problems developed in [15].
+We show that the first algorithm (Adaptive MAC) algorithm achieves an $\mathcal{O}(\sqrt{\log(T)/T})$ performance gap over any interval of $T$ consecutive time slots during which the (unknown) success probabilities remain fixed.
+This regret bound matches, up to logarithmic factors, the minimax lower bound for non-stationary multi-armed bandits established in [26].
+Moreover, if the success probabilities change at some time $T_{0}$, the performance guarantees over the interval $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$ are independent of the system behavior prior to $T_{0}$, even though the algorithm does not know the change point $T_{0}$.
 Hence, the algorithm is adaptive.
-This is possible thanks to the importance sampling technique in [29], which we use to estimate the true link failure probabilities in each time slot.
-This technique uses feedback only from the previous time slot.
-The second algorithm is based on upper confidence bound (UCB) techniques and gets within $\mathcal{O}(\sqrt{\log(T)/T})$ of optimality within a finite time horizon of $T$ time slots.
-Although this algorithm has faster convergence compared to the first algorithm, it is not adaptive due to the UCB-based estimation.
-However, the channel-user assignment of this algorithm has a max-weight structure and can be solved using the well-known Hungarian algorithm [49].
-This is in contrast to the first algorithm, where the assignment is sampled from a distribution in each time slot, which may be less desirable in certain scenarios.
-Due to the max-weight decisions, the second algorithm can be implemented in a distributed setting in the absence of a centralized controller, given that all the users have access to feedback on which channels are successfully accessed in each time slot.
-Simulations also depict the adaptiveness of the first algorithm and the faster convergence of the second algorithm.
+Adaptive MAC achieves fast convergence but requires solving a convex optimization problem in each iteration, which is computationally expensive.
+To reduce computational complexity, we propose a second adaptive algorithm (Adaptive MAC.CF) whose iterations admit closed-form solutions.
+This approach trades off a slower convergence rate for significantly lower per-slot computational cost.
+  3. 3.
+    Special Cases: In the single-channel special case, we recover both fast convergence and closed form solutions for iterations via a simpler adaptive algorithm.
+When adaptation is not required, we also consider a UCB-based algorithm that achieves the same asymptotic convergence rate while relying only on simple max-weight-type decisions.
+The UCB algorithm highlights the advantages of adaptive algorithms in non-stationary environments.
+We further study the single-channel case with utility function $\phi(\boldsymbol{x})=min\{x_{1},\dots,x_{n}\}$, where a simple estimation-free mechanism achieves both convergence and adaptation.
+  4. 4.
+    Empirical evaluation: Through simulations, we demonstrate the fast convergence of the proposed adaptive algorithms and their ability to quickly adjust to changes in link success probabilities.
+In the conducted simulations, the Adaptive MAC.CF algorithm reduces the per-iteration computational complexity by approximately 36% compared to the Adaptive MAC algorithm.
+We also empirically compare our methods with prior work in the single-channel setting.
 I-D Notation
-[…]
-I-E Definitions
-In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also, define the sets
-$\Delta_{l,\varepsilon}=\Big{\{}\boldsymbol{p}\in\mathbb{R}^{k}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\Big{\}},\text{ where }l\in\mathbb{N},$
+For integers $a,b$, we use $[a:b]$ to denote the set of integers between $a,b$ inclusive.
+We use $[a]=[1:a]$.
+For $\boldsymbol{a}\in\mathbb{R}^{k}$, $\lVert\boldsymbol{a}\rVert=\sqrt{\sum_{i=1}^{k}a_{i}^{2}}$, $\lVert\boldsymbol{a}\rVert_{1}=\sum_{i=1}^{k}|a_{i}|$, and $\boldsymbol{a}]_{+}\in\mathbb{R}^{k}$ is the vector with $i$-th entry $\max\{a_{i},0\}$.
+We use $\boldsymbol{1}_{k}$ to denote the $k$-dimensional vector of ones.
+When the dimension is clear from context, we use $\boldsymbol{1}$ instead of $\boldsymbol{1}_{k}$.
+For vectors $\boldsymbol{a,b}\in\mathbb{R}^{k}$, $\boldsymbol{c}=\boldsymbol{a}\odot\boldsymbol{b}\in\mathbb{R}^{k}$ is defined such that $c_{i}=a_{i}b_{i}$ for all $i\in[k]$.
+For matrices $\boldsymbol{A,B}\in\mathbb{R}^{k\times l}$ we use $\lVert\boldsymbol{A}\rVert=\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{l}A_{i,j}^{2}}$, $\lVert\boldsymbol{A}\rVert_{1}=\sum_{i=1}^{k}\sum_{j=1}^{l}|A_{i,j}|$, and $\boldsymbol{C}=\boldsymbol{A}\odot\boldsymbol{B}\in\mathbb{R}^{k\times l}$ is defined such that $C_{i,j}=A_{i,j}B_{i,j}$ for all $i\in[k]$ and $j\in[l]$.
+I-E Definitions and Preliminaries
+In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also,
+$\Delta^{l}_{\varepsilon}=\left\{\boldsymbol{p}\in\mathbb{R}^{l}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\right\},\text{ where }l\in\mathbb{N},$
 $\mathcal{S}^{\text{row}}_{\varepsilon}=\Bigg\{\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}:\sum_{k=1}^{s}P_{i,k}=1,P_{i,j}\geq\varepsilon\ \forall i,j\in[s]\Bigg\},$
@@
 $\mathcal{S}^{\text{doub}}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cap\mathcal{S}^{\text{row}}_{\varepsilon},\ \mathcal{S}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cup\mathcal{S}^{\text{row}}_{\varepsilon}.$
-We also denote $\Delta_{l}=\Delta_{l,0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$.
+We also denote $\Delta^{l}=\Delta^{l}_{0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$.
+Also, let $\Delta^{l}_{+}=\Delta^{l}\cap(0,1]^{l}$, $\mathcal{S}^{\text{doub}}_{+}=\mathcal{S}^{\text{doub}}\cap(0,1]^{s\times s}$ and $\mathcal{S}_{+}=\mathcal{S}\cap(0,1]^{s\times s}$.
 We hide the dependence on $s$ in the notation for sets for clarity.
+$\beta$-approximation
+Given $\beta>0$, and matrix $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{+}$, we say matrix $\hat{P}\in\mathcal{S}^{\text{doub}}_{+}$ is a $\beta$-approximation of $\boldsymbol{P}$ if
+$e^{-\beta}\leq\frac{\hat{P}_{j,j}}{P_{i,j}}$ (2)
+hold for all $i,j\in[s]$.
+KL-divergence
+For $\boldsymbol{p}\in\Delta^{s}$ and $\boldsymbol{q}\in\Delta^{s}_{+}$, we define the Kullback–Leibler (KL) divergence as
+$D_{\mathrm{KL}}(\boldsymbol{p}\lVert\boldsymbol{q})=\sum_{i=1}^{s}p_{i}\ln\left(\frac{p_{i}}{q_{i}}\right),$ (3)
+with the convention that $0\ln(0/q)=0$ for any $q>0$. For matrices $\boldsymbol{P}\in\mathcal{S}$ and $\boldsymbol{Q}\in\mathcal{S}_{+}$, we define the (entrywise) matrix KL divergence by
+$D(\boldsymbol{P}\lVert\boldsymbol{Q})=\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{P_{i,j}}{Q_{i,j}}\right),$ (4)
+where again $0\ln(0/q)=0$ for $q>0$.
+Sampling from $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$
+We define the process of sampling a permutation matrix $\boldsymbol{Y}$ from a doubly stochastic matrix $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$.
+By the Birkhoff–von Neumann theorem [12], there exists an integer $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, where $\boldsymbol{s}\in\Delta^{r}$ is a probability vector over $[1:r]$.
+Consider the following sampling procedure: draw an index $a\in[1:r]$ according to the distribution $\boldsymbol{s}$ and set $\boldsymbol{Y}=\boldsymbol{M}^{a}$.
+Then, $\mathbb{E}\!\left\{\boldsymbol{Y}\mid\boldsymbol{P}\right\}=\boldsymbol{P}$.
 I-F Assumptions
-Before moving on to the problem, we state our main assumptions.
+We state our main assumptions.
   • A1
-    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $[0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$.
+    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$.
 Hence, $\phi$ is $\sqrt{n}B$-Lipschitz continuous.
@@
     We have access to the solution of the problem $\max_{\boldsymbol{x}\in[0,1]^{n}}[\phi(\boldsymbol{x})+\sum_{i=1}^{n}c_{i}x_{i}]$, for all $\boldsymbol{c}\in\mathbb{R}^{n}_{+}$.
-    Note: This assumption is valid for most separable functions $\phi$.
-For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem has an explicit solution.
+    Note: This is valid for most separable functions $\phi$.
+For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem in A2 has an explicit solution.
 II Problem Setup
@@
 (P1:) $\underset{\boldsymbol{Y}(1),\boldsymbol{Y}(2),\dots}{\text{max}}\lim\inf_{T\to\infty}\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$ (5)
+$\text{s.t. }\boldsymbol{Y}(t),\text{ and }\boldsymbol{S}(\tau)\text{ are independent for all }$
 […]
-$\ \ \ \ \ \ \ t,\tau\in\mathbb{N}\text{ and }\tau\geq t$ (2)
 $\ \ \ \ \ \boldsymbol{Y}(t),\text{ and }S_{i,j}(\tau)\text{ are independent for all }t,\tau\in\mathbb{N},$
-$\ \ \ \ \ \ \ (i,j)\in[n]\times[m],\tau<t,Y_{i,j}(\tau)\neq 1$ (3)
-$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{n\times m}\ \forall t\in\mathbb{N}$ (4)
-$\ \ \ \ \ \sum_{i=1}^{n}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[m]$ (5)
-$\ \ \ \ \ \sum_{j=1}^{m}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[n]$ (6)
+$\ \ \ \ \ \ \ (i,j)\in[s]\times[s],\tau<t,Y_{i,j}(\tau)\neq 1$ (7)
+$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{s\times s}\ \forall t\in\mathbb{N}$ (8)
+$\ \ \ \ \ \sum_{i=1}^{s}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[s]$ (9)
+$\ \ \ \ \ \sum_{j=1}^{s}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[s]$ (10)
 $\ \ \ \ \ X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)\forall t\in\mathbb{N},\ i\in[n],$ (11)
 where constraint (6) ensures transmission decisions do not know success/failures before they happen; (7) ensures we cannot use information that is never observed.
-Define $s=\max\{n,m\}$ and $\phi^{\text{opt}}$ as the optimal objective value of (P1).
-Lemma 1
-Consider the following problem.
+Define $\phi^{\text{opt}}$ as the optimal objective value of (P1).
+II-A Deterministic Problem
+Consider the deterministic problem given by
 (P2:) $\underset{\boldsymbol{P},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (12)
-[…]
+$\text{s.t. }\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$ (13)
 $\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (14)
 $\ \ \ \ \ \sum_{j=1}^{m}q_{i,j}P_{i,j}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (15)
-and let $\phi^{*}$ denote the optimal objective value of the above problem. Then we have $\phi^{*}=\phi^{\text{opt}}$.
+Let $\phi^{*}$ denote the optimal objective value of the above problem.
+In Lemma 1, we show that $\phi^{*}\geq\phi^{\text{opt}}$.
+This lemma will be useful in establishing the performance guarantees of our algorithms.
+Let us define $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ as the optimal solution Problem (P2).
+Note: More generally, it can be shown that $\phi^{*}=\phi^{\text{opt}}$.
+If $q_{i,j}$ were known, Problem (P2) could be solved to obtain $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$.
+The stationary policy that samples a permutation matrix from $\boldsymbol{P}^{*}$ in each time slot (see Section I-E-c) and uses it for assignment achieves an objective value of $\phi^{*}$ for Problem (P1).
+However, even when $q_{i,j}$ is unknown, Section IV establishes that our algorithm attains an objective value arbitrarily close to $\phi^{*}$, thereby implying that $\phi^{\text{opt}}\geq\phi^{*}$.
+Lemma 1
+We have $\phi^{*}\geq\phi^{\text{opt}}$, where $\phi^{\text{opt}}$ and $\phi^{*}$ are the optimal objective values of Problems (P1) and  (P2), respectively.
 Proof:
-We first prove that $\phi^{*}\geq\phi^{\text{opt}}$.
-Fix $\varepsilon>0$.
-Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (2)-(6), such that
+Fix $\varepsilon>0$. Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (6)-(10), such that
 $\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}^{*}(t)\}\right)\geq\phi^{\text{opt}}-\varepsilon,$ (16)
 where $X^{*}_{i}(t)=\sum_{j=1}^{m}Y^{*}_{i,j}(t)S_{i,j}(t)$ for all $t\in\mathbb{N}$ and $i\in[n]$. Define the matrix $\boldsymbol{\tilde{P}}^{*}\in\mathbb{R}^{s\times s}$ such that
-$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if }(i,j)\in[n]\times[m]\\ 0&\text{ otherwise. }\end{cases}$
-Let us define $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\boldsymbol{P}^{*}\geq\boldsymbol{\tilde{P}}^{*}$.
-It can be observed that such a $\boldsymbol{P}^{*}$ exists from the definition of $\boldsymbol{\tilde{P}}^{*}$ and the constraints (4)-(6).
+$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if $(i,j)\in[n]\times[m]$}\\ 0&\text{ otherwise}\end{cases}$ (17)
+Constraints (8)–(10) ensure the $s\times s$ matrix $\boldsymbol{\tilde{P}}^{*}$ has nonnegative entries with all row and column sums less than or equal to 1.
+A well known result of Von-Neumann ensures that there is a matrix $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\tilde{P}^{*}_{i,j}\leq P^{*}_{i,j}$ for all $i,j\in[s]$.
 Let $\boldsymbol{\gamma}^{*}\in[0,1]^{n}$ such that $\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}$ for $i\in[n]$.
@@
 Hence, we have $\phi^{*}\geq\phi^{\text{opt}}$ as desired.
-Now, we prove that $\phi^{*}\leq\phi^{\text{opt}}$.
-Let $\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*}$ denote the optimal solution for (P2).
-Using Birkhoff-von Neumann Theorem [50], notice that there exists $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}^{*}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, and $\boldsymbol{s}\in\Delta_{r}$.
-In each time slot, we simply sample $l_{t}\sim\boldsymbol{s}(t)$ and set $\boldsymbol{Y}(t)$ according to $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for $i\in[n],j\in[m]$.
-It can be easily seen that this policy yields an objective value of $\phi^{*}$ for (P1).
-Hence, we are done.
 ∎
-III Adaptive Algorithm
+III Intuition Behind Algorithms
+Before moving onto the algorithms, we describe the general intuition unifying all the algorithms. Consider the partial Lagrangian of problem (P2) defined as
+$f(\boldsymbol{Q},\boldsymbol{P},\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}\!\left(\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right),$ (19)
+obtained by dualizing constraint (15), where $Q_{i}\geq 0$ are the Lagrange multipliers associated with (15), and $V>0$ is a tunable parameter that scales the utility term without affecting the optimal solution of (P2).
+If the success probabilities $q_{i,j}$ were known, then for a suitably chosen parameter $V$, the iterative primal update
+$(\boldsymbol{\gamma}(t+1),\boldsymbol{P}(t))$ $=\arg\min_{\begin{subarray}{c}\boldsymbol{P}\in\mathcal{S}^{\text{doub}}\\ \boldsymbol{\gamma}\in[0,1]^{n}\end{subarray}}f(\boldsymbol{Q}(t),\boldsymbol{P},\boldsymbol{\gamma})$ (20)
+$=\arg\min_{\boldsymbol{P},\boldsymbol{\gamma}}\left\{-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\right\},$
+together with the dual update
+$Q_{i}(t+1)=\left[Q_{i}(t)+\gamma_{i}(t+1)-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right]_{+},$ (21)
+for all $i\in[n]$, guarantees
+$\phi(\bar{\boldsymbol{\gamma}}(T))$ $\geq\phi^{\text{opt}}-\frac{o(T)}{T},$ (22)
+and
+$\bar{\gamma}_{i}(T)$ $\leq\sum_{j=1}^{m}q_{i,j}\bar{P}_{i,j}(T)+\frac{o(T)}{T},\quad\forall i\in[n],$ (23)
+where the time averages are defined as $(\bar{\boldsymbol{\gamma}}(T),\bar{\boldsymbol{P}}(T))=\frac{1}{T}\sum_{t=1}^{T}(\boldsymbol{\gamma}(t),\boldsymbol{P}(t))$.
+The result follows from standard primal–dual convergence arguments [39].
+Hence, we can get arbitrarily close to the optimal solution of (P2).
+We can either find $\boldsymbol{\bar{P}}(T)$ offline using the above approach, and implement the fixed policy of sampling an assignment from $\boldsymbol{\bar{P}}(T)$ (see Section I-E-c) in each time slot, or we can implement an online procedure where we sample an assignment from $\boldsymbol{P}(t)$ in time slot $t$.
+Both of these approaches will guarantee (1) given that $q_{i,j}$ values are known.
+All of our algorithms follow the structure described above.
+However, in our scenario, we do not know the values of $q_{i,j}$.
+Hence, they have to be estimated using the bandit feedback.
+For this task, in the adaptive algorithms, we use the EXP3.S importance sampling based estimation, whereas in the non-adaptive algorithm, we use UCB-based estimation.
+For each $i\in[n]$, the sequence $Q_{i}(1),Q_{i}(2),\dots$ is referred to as a virtual queue, since the dual update (21) resembles a standard queueing recursion.
+Observe that the primal update (20) decomposes into separate decisions over $\boldsymbol{\gamma}$ and $\boldsymbol{P}$.
+This decomposition is the motivation behind the introduction of the auxiliary variable $\boldsymbol{\gamma}$.
+In particular, it enables us to decouple the matching decisions from the utility optimization.
+IV Adaptive MAC Algorithm
 Now, we develop our first algorithm.
-The idea is to find $\boldsymbol{\tilde{P}}(t)$ as a column stochastic matrix and a row stochastic matrix alternatively in odd and even slots.
-Then we obtain $\boldsymbol{P}(t)$ by approximating $\boldsymbol{\tilde{P}}(t)$ by a doubly stochastic matrix, using the rounding trick similar to the one introduced in [33], after which we sample a permutation matrix from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [50].
-To do the assignment in time slot $t-1$, we discard the last $s-n$ rows or last $s-m$ columns of the permutation matrix, depending on whether $m$ or $n$ is larger.
-1 Initialize $\boldsymbol{\tilde{P}}(1)\in\mathcal{S}^{\text{doub}}_{\varepsilon}$ and the virtual queues $\boldsymbol{Q}(1)\in[0,BV+1]^{n}$ arbitrarily.
+Similar to the description of the known $q_{i,j}$ setting in Section III, the idea is to implement an online algorithm to find a doubly stochastic matrix $\boldsymbol{P}(t)$ in each time using the success/failure feedback.
+However, this is more challenging since we do not know $q_{i,j}$, and we only receive bandit feedback on success/failures.
+In addition, we require the algorithm to adapt when the success failure probabilities change.
+To achieve this, we combine EXP3.S based importance sampling idea with Lyapunov optimization.
+In each time slot, we solve a convex optimization problem to obtain an intermediate matrix $\boldsymbol{\tilde{P}}(t)$.
+The problem cannot be solved exactly.
+Hence, we find an approximation $\boldsymbol{\hat{P}}(t)$ of $\boldsymbol{\tilde{P}}(t)$.
+We then find the doubly stochastic matrix $\boldsymbol{P}(t)$ by mixing $\boldsymbol{\hat{P}}(t)$ with the all one matrix.
+This step is vital in ensuring adaptiveness.
+We then sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [12].
+We assume that if $i>n$ or $j>m$, $Y_{i,j}(t)$ corresponds to a fictional link.
+Formally, we extend the notation by setting $S_{i,j}(t)=0$ and $q_{i,j}=0$ for $i>n$ and $j>m$.
+In this algorithm, we use the loss-based importance sampling estimator
+$\hat{S}_{i,j}(t)=1-\frac{(1-S_{i,j}(t))Y_{i,j}(t)}{P_{i,j}(t)}$ (24)
+of $q_{i,j}$. This enables fast convergence guarantees.
+In Algorithm 1, we provide the algorithm for the task. The algorithm uses parameters $V>0$, $\theta>0$, $\varepsilon\in(0,1/2]$, and $\eta>0$.
+1Initialize $\boldsymbol{P}(1)=\frac{1}{s}\boldsymbol{1}\boldsymbol{1}^{\top}\in\mathcal{S}^{\text{doub}}$ and the virtual queues $\boldsymbol{Q}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$.
 2 for each time slot $t\in\mathbb{N}$ do
-3       Set $\boldsymbol{P}(t)=\text{ROUND}(\boldsymbol{\tilde{P}}(t))$ (This function yields $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ ).
-4
-5      Using Birkhoff-von Neumann Decomposition [50], find $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}(t)=\sum_{l=1}^{r}s_{l}(t)\boldsymbol{M}^{l}$, and $\boldsymbol{s}(t)\in\Delta_{r}$.
-6       Sample $l_{t}\sim\boldsymbol{s}(t)$ and take action $\boldsymbol{Y}(t)$, where $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for all $i\in[n]$, $j\in[m]$, and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
-7       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=S_{i,j}(t)Y_{i,j}(t)/P_{i,j}(t)$ for all $i\in[n]$, and $j\in[m]$.
-8       Find $\boldsymbol{\gamma}(t+1)\in[0,1]^{n}$, and $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{Q}$ using
-$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big{]}.$
-$\boldsymbol{\tilde{P}}(t+1)=\arg\min_{\boldsymbol{P}\in\mathcal{Q}}\Big{[}-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\Big{]},$ (14)
-where $\mathcal{Q}=\mathcal{S}^{\text{col}}_{\varepsilon}$ if $t$ is even, and $\mathcal{Q}=\mathcal{S}^{\text{row}}_{\varepsilon}$ if $t$ is odd, and the divergence $D$ is defined by
-$D(\boldsymbol{X}\lVert\boldsymbol{Y})=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right).$ (15)
-for all $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}.$
+3   Sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ (see Section I-E-c), and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
+4   Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=1-\frac{(1-S_{i,j}(t))Y_{i,j}(t)}{P_{i,j}(t)}$ for all $i,j\in[s]$ where we define $S_{i,j}(t)=0$ whenever $i>n$ or $j>m$.
+5   Define
+$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big[-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big].$ (25)
+6   Find $\boldsymbol{\hat{P}}(t+1)$ as a $(\theta/s)$-approximation of $\boldsymbol{\tilde{P}}(t+1)$ where
+$\boldsymbol{\tilde{P}}(t+1)$ $=\arg\min_{\boldsymbol{P}\in\mathcal{S}^{\text{doub}}}\Big[-\sum_{i=1}^{n}\sum_{j=1}^{s}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\Big],$ (26)
+where $\beta$-approximation is defined in (2) and the divergence $D(\cdot\lVert\cdot)$ is defined in (4).
+7   Mix
+$\boldsymbol{P}(t+1)=(1-\varepsilon)\boldsymbol{\hat{P}}(t+1)+\frac{\varepsilon}{s}\boldsymbol{1}\boldsymbol{1}^{\top}$ (27)
+8
 9  Update the virtual queues
@@
 10
-Algorithm 1 Adaptive MAC
-Below, we introduce the ROUND function, a technique adapted from the one introduced in [33] to approximate a nonnegative matrix by a matrix in the transport polytope.
-The technique can be readily extended to the Birkhoff polytope using the same algorithm.
+Algorithm 1 Adaptive MAC (Parameters $V>0$, $\theta>0$, $\varepsilon\in(0,1/2]$, and $\eta>0$)
+The following Theorem establishes the performance bound of Algorithm 1.
+Theorem 1
+Consider $T,T_{0}\in\mathbb{N}$ and suppose that the success probabilities $q_{i,j}$ remain fixed on the interval $T_{0},\,T_{0}+T-1]$.
+Then, independently of their behavior outside this interval, for any algorithm parameters $\varepsilon\in(0,1/2]$, $\eta>0$, $\theta>0$, and $V>0$, Algorithm 1 yields
+$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$
+$\leq\frac{n}{V}+\frac{2\varepsilon s}{\eta V}+\frac{\theta}{V\eta}+\frac{\eta ns(BV+1)^{2}}{V}+\frac{s}{\eta VT}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
+where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1.
+In particular, choosing $\eta=\Theta(\sqrt{\log(T)}/T)$, $V=\Theta(\sqrt{T})$, $\varepsilon=\Theta(1/T)$, and $\theta=\Theta(1/T)$, we have
+$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\sqrt{\frac{\log(T)}{T}}\right).$
+IV-A Discussion of Algorithm 1
+In this section, we connect Algorithm 1 with the Intuition described in Section III. Define the history $\mathcal{H}(t)$ as the sigma algebra generated by
+$\mathcal{H}(t)=\{\boldsymbol{Y}(\tau),\;\boldsymbol{Y}(\tau)\odot\boldsymbol{S}(\tau):\ \tau\in[t-1]\}.$ (29)
+Using the importance-sampling estimator $\hat{S}_{i,j}(t)$ (see line 5 of Algorithm 1), we can form unbiased estimates of the quantities $q_{i,j}$ appearing in the primal objective (20).
+In particular, for any $\mathcal{H}(t)$-measurable $\boldsymbol{P}\in\mathcal{S}^{\mathrm{doub}}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{Q}\geq 0$, the random variable
+$-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}\!\left[\gamma_{i}-\sum_{j=1}^{m}\hat{S}_{i,j}(t)P_{i,j}\right]$ (30)
+is an unbiased estimator of $f(\boldsymbol{Q},\boldsymbol{P},\boldsymbol{\gamma})$ defined in (19).
+However, we cannot directly replace the primal update (20) at time $t$ by the estimator (30), since $\hat{S}_{i,j}(t)$ depends on the decision $\boldsymbol{P}(t)$ over which we are optimizing (see line 5 of Algorithm 1).
+To enable an update of the choice probabilities, we instead introduce a new decision variable $\boldsymbol{\tilde{P}}(t+1)$ and compute $(\boldsymbol{\gamma}(t+1),\boldsymbol{\tilde{P}}(t+1))$ by minimizing the surrogate objective
+$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}Q_{i}(t)\!\left[\gamma_{i}(t+1)-\sum_{j=1}^{m}\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right]+\frac{D\big(\boldsymbol{\tilde{P}(t+1)}\lVert\boldsymbol{P}(t)\big)}{\eta}.$ (31)
+Since the optimization problem in (31) does not admit a closed-form solution, we compute $\boldsymbol{\hat{P}}(t+1)$ as an approximation to $\boldsymbol{\tilde{P}}(t+1)$.
+Although the first two terms of (31) do not form an unbiased estimator of $f(\boldsymbol{Q}(t),\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1))$ (since $\boldsymbol{\tilde{P}}(t+1)$ is not $\mathcal{H}(t)$-measurable), the divergence penalty forces $\boldsymbol{\tilde{P}}(t+1)$ to remain close to the $\mathcal{H}(t)$-measurable point $\boldsymbol{P}(t)$.
+Intuitively, this ensures that the surrogate objective remains close to the true primal objective evaluated at $\boldsymbol{P}(t)$, while still allowing the policy to evolve over time.
+We mix $\boldsymbol{\hat{P}}(t+1)$, the approximation of $\boldsymbol{\tilde{P}}(t+1)$, with the uniform doubly stochastic matrix $\frac{1}{s}\boldsymbol{1}\boldsymbol{1}^{\top}$ to obtain $\boldsymbol{P}(t+1)$ in line 8 of Algorithm 1.
+This guarantees persistent exploration and ensures that the algorithm remains adaptive.
+Finally, note that in the virtual queue update (21), we replace the last term with the unbiased estimator $\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$. In particular,
+$\mathbb{E}\left\{\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)\,\Big|\,\mathcal{H}(t)\right\}=\sum_{j=1}^{m}P_{i,j}(t)q_{i,j}.$ (32)
+This estimator can be computed under bandit feedback, since it only requires observing $S_{i,j}(t)$ for indices with $Y_{i,j}(t)=1$.
+IV-B Problems (25)-(26)
+Finding $\boldsymbol{\gamma}(t+1)$ (Problem (25)) is possible due to Assumption A2.
+A standard approach for computing $\boldsymbol{\hat{P}}(t+1)$ (Problem (26)) is to apply Sinkhorn procedure, as discussed in Appendix B.
+Each Sinkhorn iteration has computational complexity $\mathcal{O}(s^{2})$.
+Consequently, the overall complexity of computing $\boldsymbol{\hat{P}}(t+1)$ is $\mathcal{O}(s^{2}T^{\text{inner}})$, where $T^{\text{inner}}$ denotes the number of Sinkhorn iterations performed.
+While Sinkhorn’s procedure exhibits linear convergence in our problem, the exact convergence rate depends on the matrices $\boldsymbol{P}(t)$, $\boldsymbol{\hat{S}}(t)$, and the vector $\boldsymbol{Q}(t)$.
+To mitigate this dependence on the inner convergence behavior, we propose in Section V the Adaptive MAC.CF algorithm, in which each time slot has computational complexity $\mathcal{O}(s^{2})$.
+The rest of this section focuses on proving Theorem 1.
+IV-C Preliminary Lemmas
+We state the preliminary lemmas used in the proof of Theorem 1. We begin with a standard result on minimizing strongly convex functions (see, e.g., [49]).
+Lemma 2
+Consider $\beta\in[0,1/s]$ and $\alpha>0$. Let
+  1. 1.
+    $\mathcal{Q}\in\{\mathcal{S}^{\text{doub}}_{\beta},\mathcal{S}^{\text{row}}_{\beta},\mathcal{S}^{\text{col}}_{\beta},\mathcal{S}_{\beta}\}$, $g:\mathbb{R}_{+}^{s\times s}\to\mathbb{R}$ be a convex function, $\boldsymbol{Y}\in\mathcal{S}_{+}$, and $\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}\left[g(\boldsymbol{X})+\alpha D(\boldsymbol{X}\lVert\boldsymbol{Y})\right]$.
+We have for all $\boldsymbol{Z}\in\mathcal{Q}$,
+    $g(\boldsymbol{X}^{*})+\alpha D(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}).$
+  2. 2.
+    $\mathcal{Q}=\Delta^{s}_{\varepsilon}$, $g:\mathbb{R}_{+}^{s}\to\mathbb{R}$ be a convex function, $\boldsymbol{Y}\in\Delta^{s}_{+}$, and $\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}[g(\boldsymbol{X})+\alpha D_{\text{KL}}(\boldsymbol{X}\lVert\boldsymbol{Y})]$.
+We have for all $\boldsymbol{Z}\in\mathcal{Q}$,
+    $g(\boldsymbol{X}^{*})+\alpha D_{\text{KL}}(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D_{\text{KL}}(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D_{\text{KL}}(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}).$
+Lemma 3
+We have for $\boldsymbol{X}\in\mathcal{S}$, $\boldsymbol{Y}\in\mathcal{S}_{+}$,
+$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert^{2},$ (33)
+Proof:
+Notice that
+$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\sum_{i=1}^{s}\frac{1}{2}\left(\sum_{j=1}^{s}|X_{i,j}-Y_{i,j}|\right)^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}$
+where the first inequality follows from the Pinsker’s inequality and the last inequality follows from the fact that for $a_{1},a_{2},\dots,a_{s}\in\mathbb{R}$, $\sum_{i=1}^{s}a_{i}^{2}\geq\left(\sum_{i=1}^{s}a_{i}\right)^{2}/s$.
+∎
+Lemma 4
+We have $D(\boldsymbol{X}\lVert\boldsymbol{Y})\leq s\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{X}\in\mathcal{S}$, where $\boldsymbol{Y}\in\mathcal{S}_{\varepsilon}$.
+Proof:
+Notice that
+$D(\boldsymbol{X}\lVert\boldsymbol{Y})$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right)$
+$\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{1}{\varepsilon}\right)=s\log\left(\frac{1}{\varepsilon}\right),$
+where (a) follows since $X_{i,j}\leq 1$, $Y_{i,j}\geq\varepsilon$, and log is a non-decreasing function. ∎
+Lemma 5
+Consider $\boldsymbol{x}\in\mathbb{R}^{s}$ and $\boldsymbol{p}\in\Delta^{s}$, $\boldsymbol{q}\in\Delta^{s}_{+}$. We have
+$\sum_{i=1}^{s}p_{i}x_{i}\leq D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{q})+\ln\left(\sum_{i=1}^{s}q_{i}e^{x_{i}}\right)$ (34)
+Proof:
+Let $A=\sum_{j=1}^{s}q_{j}e^{x_{j}}$. Define the vector $\boldsymbol{r}\in\Delta_{+}$ by $r_{i}=q_{i}e^{x_{i}}/A$. Notice that
+$D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{r})$ $=\sum_{i=1}^{s}p_{i}\ln\left(\frac{p_{i}}{r_{i}}\right)=\sum_{i=1}^{s}p_{i}\ln\left(\frac{p_{i}A}{q_{i}e^{x_{i}}}\right)=D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{q})+\ln(A)-\sum_{i=1}^{s}p_{i}x_{i}$ (35)
+Using $D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{r})\geq 0$ and rearranging, we are done. ∎
+Now, we establish the following lemma, which will be useful in establishing a deterministic bound on $Q_{i}(t)$.
+Lemma 6
+Consider the three vector sequences $\{\boldsymbol{A}(t)\in[0,1]^{n}:t\in\mathbb{N}\},\{\boldsymbol{G}(t)\in[0,1]^{n}:t\in\mathbb{N}\},\{\boldsymbol{F}(t)\in[0,\infty)^{n}:t\in\mathbb{N}\}$, where $\{\boldsymbol{A}(t):t\in\mathbb{N}\}$ is an arbitrary sequence in $0,1]^{n}$, and
+$\boldsymbol{F}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$
+$\boldsymbol{G}(t+1)=\arg\min_{\boldsymbol{g}\in[0,1]^{n}}\left[-V\phi(\boldsymbol{g})+\sum_{i=1}^{n}F_{i}(t)g_{i}\right],\ \forall t\in\mathbb{N},$
+$\boldsymbol{F}(t+1)=[\boldsymbol{F}(t)+\boldsymbol{G}(t+1)-\boldsymbol{A}(t)]_{+}\ \forall t\in\mathbb{N}$ (36)
+We have for all $t\in\mathbb{N}$ and $i\in[n]$, $F_{i}(t)\leq BV+1$, where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1.
+Proof:
+Fix $t\in\mathbb{N}$. We first prove the following claim.
+Claim: If $F_{k}(t)>BV$, for $k\in[n]$, then $G_{k}(t+1)=0$.
+Proof:
+We prove the contrapositive. In particular, assume that $G_{k}(t+1)>0$. We prove $F_{k}(t)\leq BV$. Define the function $f:[0,1]^{n}\to\mathbb{R}$ by
+$f(\boldsymbol{g})=-V\phi(\boldsymbol{g})+\sum_{i=1}^{n}g_{i}F_{i}(t)$ (37)
+Notice that due to the definition of $\boldsymbol{G}(t+1)$ we have
+$f(\boldsymbol{G}(t+1))\leq f(\boldsymbol{g})$ (38)
+for any $\boldsymbol{g}\in[0,1]^{n}$. Using the subgradient inequality for the convex function $f$, we have for any $\boldsymbol{g}\in[0,1]^{n}$
+$f^{{}^{\prime}}(\boldsymbol{g})[\boldsymbol{G}(t+1)-\boldsymbol{g}]\leq f(\boldsymbol{G}(t+1))-f(\boldsymbol{g})\leq 0$ (39)
+where the last inequality follows from (38). Let us pick $\boldsymbol{g}$ as
+$g_{i}=\begin{cases}0&\text{ if }i=k\\ G_{i}(t+1)&\text{ otherwise}\end{cases}$ (40)
+Using (39) with the above $\boldsymbol{g}$, we have
+$0$ $\geq f^{{}^{\prime}}(\boldsymbol{g})[\boldsymbol{G}(t+1)-\boldsymbol{g}]=\sum_{i=1}^{n}[-V\phi_{i}^{{}^{\prime}}(\boldsymbol{g})+F_{i}(t)](G_{i}(t+1)-g_{i})$
+$=[-V\phi_{k}^{{}^{\prime}}(\boldsymbol{g})+F_{i}(t)]G_{k}(t+1)$ (41)
+where for the last equality, we have used the definition of $\boldsymbol{g}$ in (40).
+Notice that since by assumption $G_{k}(t+1)>0$, the above translates to $V\phi_{k}^{{}^{\prime}}(\boldsymbol{g})+F_{i}(t)\leq 0$.
+Since we have $|\phi_{k}^{{}^{\prime}}(\boldsymbol{g})|\leq B$ (Assumption A1), we have $F_{i}(t)\leq V\phi_{k}^{{}^{\prime}}(\boldsymbol{g})\leq BV$ as desired.
+∎
+Now, we move on to the proof of Lemma 6.
+Fix $i\in[n]$.
+We use induction to prove that $F_{i}(t)\leq BV+1$ for all $t\in\mathbb{N}$.
+Notice that the result trivially follows for $t=1$ ($\boldsymbol{F}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$).
+Assume $F_{i}(t)\leq BV+1$.
+We establish $F_{i}(t+1)\leq BV+1$.
+We consider two cases.
+Case 1: $F_{i}(t)=BV+1$: Notice that from the previous claim we have $G_{i}(t+1)=0$. Hence, we have from the definition of $F(t+1)$ in (6)
+$F_{i}(t+1)$ $=[F_{i}(t)+G_{i}(t+1)-A_{i}(t)]_{+}\leq_{(a)}[F_{i}(t)]_{+}=BV+1,$
+where (a) follows since $G_{i}(t+1)=0$, $A_{i}(t)\geq 0$, and $\cdot$$_{+}$ is a non-decreasing function, and the last inequality follows from $F_{i}(t)=BV+1$.
+Case 2: $F_{i}(t)\leq BV$: We have from (6),
+$F_{i}(t+1)$ $=[F_{i}(t)+G_{i}(t+1)-A_{i}(t)]_{+}\leq_{(a)}[F_{i}(t)+1]_{+}\leq BV+1,$
+where (a) follows since $G_{i}(t+1)\leq 1$, $A_{i}(t)\geq 0$, and $\cdot$$_{+}$ is a non-decreasing function, and the last inequality follows from $F_{i}(t)\leq BV$ and $\cdot$$_{+}$ is non-decreasing.
+∎
+IV-D Analysis of the Algorithm
+Now, we establish the core lemmas required to prove Theorem 1.
+In this section, we consider the interval $T_{0},\,T_{0}+T-1]$, where $T,T_{0}\in\mathbb{N}$.
+We assume only that the success probabilities $q_{i,j}$ remain fixed over this interval and we impose no assumptions on their behavior outside of it.
+Before proving Theorem 1, we establish a deterministic bound $Q_{i}(t)$, which is direct consequence of Lemma 6.
+Corollary 1.1
+We have for $t\in[T_{0},\,T_{0}+T-1]$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$.
+Note: This sample path result does not require fixed $q_{i,j}$ probabilities. It only requires $X_{i}(t)\in[0,1]$, and $\gamma_{i}(t)\in[0,1]$.
+Now we state Lemma 7 that controls the error introduced by replacing the true link success probabilities with importance-sampling estimates.
+Lemma 7
+We have for all $t\in[T_{0},\,T_{0}+T-1]$,
+$-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{P_{i,j}(t)Q_{i}(t)\}$
+$\leq-\sum_{i=1}^{n}\sum_{j=1}^{s}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\}+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))+\eta(BV+1)^{2}ns$ (42)
+where $\hat{S}_{i,j}(t)$ is defined in line 5 of Algorithm 1.
+Proof:
+Consider $\boldsymbol{x}^{i},\boldsymbol{p}^{i},\boldsymbol{q}^{i}\in\mathbb{R}^{s}$, where
+$x^{i}_{j}$ $=-\eta Q_{i}(t)\bigl(1-\hat{S}_{i,j}(t)\bigr),p^{i}_{j}=\tilde{P}_{i,j}(t+1),q^{i}_{j}=P_{i,j}(t).$
+Using Lemma 5 on $\boldsymbol{x}^{i},\boldsymbol{p}^{i},\boldsymbol{q}^{i}$ we get
+$-\sum_{j=1}^{s}Q_{i}(t)(1-\hat{S}_{i,j}(t))\tilde{P}_{i,j}(t+1)$ (43)
+$\leq_{(a)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{\tilde{P}}_{i}(t+1)\lVert\boldsymbol{P}_{i}(t))+\ln\left(\sum_{j=1}^{s}{P}_{i,j}(t)e^{-\eta Q_{i}(t)(1-\hat{S}_{i,j}(t))}\right)\Bigg]$ (44)
+where $\boldsymbol{\tilde{P}}_{i}(t+1)$ and $\boldsymbol{P}_{i}(t)$ denote the $i$-th rows of $\boldsymbol{\tilde{P}}(t+1)$ and $\boldsymbol{P}_{i}(t)$, respectively. Notice that
+$D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))$ $=\sum_{i=1}^{s}D_{\text{KL}}(\boldsymbol{\tilde{P}}_{i}(t+1)\lVert\boldsymbol{P}_{i}(t))\geq\sum_{i=1}^{n}D_{\text{KL}}(\boldsymbol{\tilde{P}}_{i}(t+1)\lVert\boldsymbol{P}_{i}(t))$ (45)
+Summing (43) over $i\in[n]$ and using the above inequality, we get
+$-\sum_{i=1}^{n}\sum_{j=1}^{s}Q_{i}(t)(1-\hat{S}_{i,j}(t))\tilde{P}_{i,j}(t+1)$ (46)
+$\leq_{(a)}\frac{1}{\eta}\Bigg[D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))+\sum_{i=1}^{n}\ln\left(\sum_{j=1}^{s}{P}_{i,j}(t)e^{-\eta Q_{i}(t)(1-\hat{S}_{i,j}(t))}\right)\Bigg]$
+$\leq_{(b)}\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))$
+[…]
+$\leq\frac{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))}{\eta}$
+[…]
+$\leq_{(c)}\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))$
+$\ \ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)Q_{i}(t)(1-\hat{S}_{i,j}(t))+\eta\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)\left(Q_{i}(t)(1-\hat{S}_{i,j}(t))\right)^{2}$
+$\leq_{(d)}\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))$
+$\ \ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)Q_{i}(t)(1-\hat{S}_{i,j}(t))+\eta(BV+1)^{2}\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)(1-\hat{S}_{i,j}(t))^{2}$
+$=_{(e)}\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))$
+$\ \ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)Q_{i}(t)(1-\hat{S}_{i,j}(t))+\eta(BV+1)^{2}\sum_{i=1}^{n}\sum_{j=1}^{s}(1-S_{i,j}(t))Y_{i,j}(t)(1-\hat{S}_{i,j}(t))$
+$\leq\frac{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))}{\eta}-\sum_{i=1}^{n}\sum_{j=1}^{s}P_{i,j}(t)Q_{i}(t)(1-\hat{S}_{i,j}(t))+\eta(BV+1)^{2}\sum_{i=1}^{n}\sum_{j=1}^{s}(1-\hat{S}_{i,j}(t))$
+where (a) follows from Lemma 5, (b) follows from $e^{-x}\leq 1-x+x^{2}$ for $x\geq-1$ (notice that $Q_{i}(t)\geq 0$, and $(1-\hat{S}_{i,j}(t))\geq 0$ by definition), (c) follows from $\ln(1+x)\leq x$, and (d) follows from Corollary 1.1, (e) follows from the definition of $\hat{S}_{i,j}(t)$, and the last inequality follows from $(1-\hat{S}_{i,j}(t))\geq 0$, and $(1-S_{i,j}(t)Y_{i,j}(t))\in\{0,1\}$.
+Taking the expected value conditioned on $\mathcal{H}(t)$, followed by taking unconditional expectations, we have
+$-\sum_{i=1}^{n}\sum_{j=1}^{s}\mathbb{E}\{Q_{i}(t)(1-\hat{S}_{i,j}(t))\tilde{P}_{i,j}(t+1)\}$
+$\leq\frac{\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))\}}{\eta}-\sum_{i=1}^{n}\sum_{j=1}^{s}\mathbb{E}\{(1-q_{i})P_{i,j}(t)Q_{i}(t)\}+\eta(BV+1)^{2}\sum_{i=1}^{n}\sum_{j=1}^{s}(1-q_{i,j})$
+$\leq\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))\}-\sum_{i=1}^{n}\sum_{j=1}^{s}\mathbb{E}\{(1-q_{i})P_{i,j}(t)Q_{i}(t)\}+\eta(BV+1)^{2}ns$
+Noticing that $\sum_{j=1}^{s}\tilde{P}_{i,j}(t+1)=\sum_{j=1}^{s}P_{i,j}(t)=1$ and rearranging we have the result. ∎
+For $t\in[T_{0},\,T_{0}+T-1]$, define the drift $\Delta(t)$ as
+$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}.$
+Lemma 8 provides an upper bound on the drift.
+Lemma 10 shows that decisions (25) and (26) minimize an upper bound on the drift-plus-penalty term $\Delta(t)-V\phi(\boldsymbol{\gamma}(t+1))$.
+Intuitively, the penalty term $V\phi(\boldsymbol{\gamma}(t+1))$ drives the algorithm toward utility maximization and enforces constraint (22), while the drift term controls the growth of the virtual queues.
+Keeping the drift small promotes stability of the virtual queues, which in turn enforces the long-term constraint (23).
+The parameter $V$ balances these two effects: larger $V$ places more emphasis on utility maximization at the cost of larger virtual queues.
+Lemma 8
+We have for all $t\in[T_{0},\,T_{0}+T-1]$,
+$\Delta(t)$ $\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
+Proof:
+Notice that from the queuing equation (28), we have for all $i\in[n]$,
+$Q^{2}_{i}(t+1)$ $\leq\left(Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\right)^{2}$
+$\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t+1)+X^{2}_{i}(t)+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right]$
+$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right],$
+where for the last inequality, we have used $X_{i}(t)\in\{0,1\}$. Summing the above for $i\in[n]$, we have
+$\lVert\boldsymbol{Q}(t+1)\rVert^{2}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right].$
+Taking the expectations conditioned $\mathcal{H}(t)$:
+$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\mathbb{E}\{\gamma_{i}(t+1)|\mathcal{H}(t)\}-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right].$
+Taking expectations and performing simple algebraic manipulations, we have the result. ∎
+We introduce the following lemma that combines $\boldsymbol{\tilde{P}}(t)$, and $\boldsymbol{P}(t)$, the actual and approximate solutions of (26).
+Lemma 9
+We have that for any $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$
+$D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\leq 2\varepsilon s+\theta+D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))$ (47)
+Proof:
+Notice that, for any $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$,
+$D(\boldsymbol{P}\lVert\boldsymbol{\hat{P}}(t))=\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{P_{i,j}}{\hat{P}_{i,j}(t)}\right)\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{e^{\theta/s}P_{i,j}}{\tilde{P}_{i,j}(t)}\right)=\theta+D_{KL}(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))$
+where (a) follows since $\ln$ is non-decreasing and since $\boldsymbol{\hat{P}}(t)$ is a $(\theta/s)$-approximation of $\boldsymbol{\tilde{P}}(t)$ (See line 7 of Algorithm 1). Next,
+$D(\boldsymbol{P}\lVert\boldsymbol{P}(t))=\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{P_{i,j}}{P_{i,j}(t)}\right)\leq\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{P_{i,j}}{(1-\varepsilon)\hat{P}_{i,j}(t)}\right)$
+$=s\ln\left(\frac{1}{1-\varepsilon}\right)+D_{KL}(\boldsymbol{P}\lVert\boldsymbol{\hat{P}}(t))\leq 2\varepsilon s+D_{KL}(\boldsymbol{P}\lVert\boldsymbol{\hat{P}}(t))$
+where for the last inequality we have used $\ln\left(\frac{1}{1-\varepsilon}\right)\leq 2\varepsilon$ whenever $\varepsilon\in[0,1/2]$. Combining the two inequalities, we are done. ∎
+Now, we introduce the following lemma that will be useful in deriving the final bounds.
+Lemma 10
+We have for any $T,T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$,
+$VT\phi(\boldsymbol{\gamma})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+nT+\frac{2\varepsilon sT}{\eta}$
+$\ \ \ \ +\frac{\theta T}{\eta}+\eta ns(BV+1)^{2}T+\frac{s}{\eta}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}.$
+Proof:
+Adding $V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
+$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
+$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{s}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\}$
+$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))\}+\eta ns(BV+1)^{2},$ (48)
+where (a) follows by Lemma 7.
+Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (26) (see Algorithm 1) with Lemma 2-1, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$,
+$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))-\sum_{i=1}^{n}\sum_{j=1}^{s}Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)$
+$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{s}\hat{S}_{i,j}(t)P_{i,j}\right]+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{P}(t))-\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))$
+$\leq_{(a)}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{m}\hat{S}_{i,j}(t)P_{i,j}\right]+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{P}(t))-\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{P}(t+1))+\frac{2\varepsilon s}{\eta}$
+$\ \ \ \ +\frac{\theta}{\eta}$ (49)
+where (a) follows from Lemma 9. Taking expectations of the above, we have
+$-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{P}(t))\}$
+$\ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\}$
+$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{s}\hat{S}_{i,j}(t)P_{i,j}\right]\right\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\}$
+$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t+1))\}+2\varepsilon s+\frac{\theta}{\eta}$
+$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\}-\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t+1))\}$
+$\ \ \ \ +\frac{2\varepsilon s}{\eta}+\frac{\theta}{\eta}$
+Substituting the above in (10), we have
+$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq n-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]$
+$\ \ +\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\}-\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{P}(t+1))\}+\frac{2\varepsilon s}{\eta}+\frac{\theta}{\eta}+\eta ns(BV+1)^{2}$ (50)
+Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T+T_{0}-1\}$, we have
+$\frac{1}{2}$ $\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+nT+\frac{2\varepsilon sT}{\eta}+\frac{\theta T}{\eta}$
+$\ \ \ \ +\eta ns(BV+1)^{2}T+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{P}(T_{0}))\right\}-\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{P}(T+T_{0}))\right\}$
+$\leq-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i,j}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+nT+\frac{2\varepsilon sT}{\eta}+\frac{\theta T}{\eta}$
+$\ \ \ +\eta ns(BV+1)^{2}T+\frac{s}{\eta}\log\left(\frac{s}{\varepsilon}\right),$ (51)
+where for the last inequality, we have used Lemma 4, and the fact that $\boldsymbol{P}(T_{0})\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$.
+Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Corollary 1.1, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done.
+∎
+Next, we have the following lemma that combines iterates $\boldsymbol{\gamma}(T_{0}+1),\boldsymbol{\gamma}(T_{0}+2)\dots,\boldsymbol{\gamma}(T+T_{0})$, and $\boldsymbol{X}(T_{0}),\boldsymbol{X}(T_{0}+1)$ $\dots,\boldsymbol{X}(T+T_{0}-1)$.
+Lemma 11
+We have for any $T,T_{0}\in\mathbb{N}$,
+$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T}.$
+Proof:
+Notice that
+$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)=\phi\left(\sum_{t=T_{0}}^{T+T_{0}-1}\frac{\boldsymbol{X}(t)}{T}+\sum_{t=T_{0}}^{T+T_{0}-1}\frac{\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)}{T}\right)$
+$\leq_{(a)}\phi\left(\sum_{t=T_{0}}^{T+T_{0}-1}\frac{\boldsymbol{X}(t)}{T}+\sum_{t=T_{0}}^{T+T_{0}-1}\frac{[\boldsymbol{Q}(t+1)-\boldsymbol{Q}(t)]}{T}\right)$
+$\leq_{(b)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{\boldsymbol{Q}(T+T_{0})}{T}\right)$
+$\leq_{(c)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T},$
+where (a) follows from the entrywise nondecreasing property of $\phi$, and the queuing equation (28), for (b) we use $\boldsymbol{Q}(T_{0})\geq 0$ and the entrywise nondecreasing property of $\phi$, and (c) follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.
+[^2] Assumption A1 ensures that $\phi$ can be extended to an entrywise nondecreasing function that is $\sqrt{nB}$-Lipschitz continuous over the domain $0,\infty)^{n}$.
+Now, we are ready to establish Theorem 1.
+Proof:
+Substituting $\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*}$ in Lemma 10, where $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of (P2) defined in Lemma 1 we have
+$VT\phi(\boldsymbol{\gamma}^{*})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma^{*}_{i}-\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+nT+\frac{2\varepsilon sT}{\eta}+\frac{\theta T}{\eta}+\eta ns(BV+1)^{2}T$
+$\ \ \ \ +\frac{s}{\eta}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
+$\leq nT+\frac{2\varepsilon sT}{\eta}+\frac{\theta T}{\eta}+\eta ns(BV+1)^{2}T+\frac{s}{\eta}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (52)
+where the last inequality follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2).
+Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain
+$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
+$\leq_{(a)}\frac{n}{V}+\frac{2\varepsilon s}{\eta V}+\frac{\theta}{V\eta}+\frac{\eta ns(BV+1)^{2}}{V}+\frac{s}{\eta VT}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$ (53)
+Combining with Lemma 11, we have
+$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
+$\leq_{(a)}\frac{n}{V}+\frac{2\varepsilon s}{\eta V}+\frac{\theta}{V\eta}+\frac{\eta ns(BV+1)^{2}}{V}+\frac{s}{\eta VT}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
+$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
+$\leq_{(b)}\frac{n}{V}+\frac{2\varepsilon s}{\eta V}+\frac{\theta}{V\eta}+\frac{\eta ns(BV+1)^{2}}{V}+\frac{s}{\eta VT}\log\left(\frac{s}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
+where (a) follows due to Lemma 11, and (b) follows due to Corollary 1.1.
+Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{*}\geq\phi^{\text{opt}}$.
+Using Jensen’s inequality, we are done.
+∎
+V Adaptive MAC.CF Algorithm
+Now, we develop our second algorithm.
+The main difficulty in implementing the Algorithm 2 is the complexity of solving the inner problem (26).
+In particular, the computational complexity of solving the above problem depends heavily on the structure of the matrices $\boldsymbol{P}(t),\boldsymbol{Q}(t),\boldsymbol{\hat{S}}(t)$ [24].
+The goal of this section is to develop an adaptive algorithm that avoids the above computational complexities.
+The resulting algorithm admits efficient inner iterations, at the cost of a slower convergence rate.
+The algorithm has the structure explained in Section III.
+However, instead of finding a doubly stochastic matrix by solving an inner problem, we find $\boldsymbol{\tilde{P}}(t)$ as a column stochastic matrix and a row stochastic matrix alternatively in odd and even slots.
+Unlike the previous algorithm, the above step can be solved explicitly.
+Then we obtain $\boldsymbol{P}(t)$ by approximating $\boldsymbol{\tilde{P}}(t)$ by a doubly stochastic matrix, using the rounding trick introduced in Sections V-A, after which we sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [12].
+In Algorithm 2, we provide the algorithm for the task.
+The algorithm uses parameters $\eta>0,V>0,\varepsilon\in(0,1/s]$
+1Initialize $\boldsymbol{\tilde{P}}(1)=\frac{1}{s}\boldsymbol{1}\boldsymbol{1}^{\top}\in\mathcal{S}^{\text{doub}}$ and the virtual queues $\boldsymbol{Q}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$.
+2 for each time slot $t\in\mathbb{N}$ do
+3   Set $\boldsymbol{P}(t)=\text{ROUND}(\boldsymbol{\tilde{P}}(t))$ (See Section V-A. This function yields $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ ).
+4
+5  Sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ (see Section I-E-c), and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
+6   Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=S_{i,j}(t)Y_{i,j}(t)/P_{i,j}(t)$ for all $i\in[n]$, and $j\in[m]$.
+7   Define
+$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big[-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big]$ (54)
+8   Find $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{Q}$ using
+$\boldsymbol{\tilde{P}}(t+1)$ $=\arg\min_{\boldsymbol{P}\in\mathcal{Q}}\Big[-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\Big],$ (55)
+where $\mathcal{Q}=\mathcal{S}^{\text{col}}_{\varepsilon}$ if $t$ is even, and $\mathcal{Q}=\mathcal{S}^{\text{row}}_{\varepsilon}$ if $t$ is odd.
+9   Update the virtual queues
+$\boldsymbol{Q}(t+1)=\left[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right]_{+},$ (56)
+where $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for $i\in[n]$.
+10
+Algorithm 2 Adaptive MAC.CF (Parameters $\eta>0,V>0,\varepsilon\in(0,1/s]$)
+In this algorithm, we employ the utility-based importance sampling estimator $\hat{S}_{i,j}(t)=\frac{S_{i,j}(t)Y_{i,j}(t)}{P_{i,j}(t)}$ of $q_{i,j}$.
+Moreover, the algorithm does not require mixing with the uniform distribution, since the inner problem (26) can be solved explicitly while enforcing the constraint that each entry is at least $\varepsilon$.
+The solution is discussed in Appendix D.
+Finding $\boldsymbol{\gamma}(t+1)$ is the same procedure as in Algorithm 1.
+The following Theorem establishes the perforance bound of the Algorithm.
+Theorem 2
+Consider $T,T_{0}\in\mathbb{N}$ and suppose that the success probabilities $q_{i,j}$ remain fixed on the interval $T_{0},\,T_{0}+T-1]$.
+Then, independently of their behavior outside this interval, for any algorithm parameters $\varepsilon\in(0,1/s]$, $\eta>0$, and $V>0$, Algorithm 2 yields
+$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$
+$\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}.$
+In particular, choosing $\eta=\Theta(\log^{2/3}(T)/T)$, $\varepsilon=\Theta((\log(T)/T)^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
+$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\sqrt[3]{\frac{\log(T)}{T}}\right).$
+The basic intuition behind the algorithm is similar to that of Algorithm 1.
+The key difference lies in the separate treatment of even and odd iterations.
+The divergence term in (55) ensures that $\boldsymbol{\tilde{P}}(t+1)$ remains close to $\boldsymbol{\tilde{P}}(t)$.
+Importantly, one of these matrices is row-stochastic, whereas the other is column-stochastic.
+Therefore, although we do not enforce double stochasticity at each iteration, the divergence term guarantees that a selected row-stochastic matrix remains close to being column-stochastic, and vice versa.
+We first describe the ROUND function used in line 3 of the algorithm.
+V-A ROUND function
+We introduce ROUND, a technique adapted from the one introduced in [5] to approximate a nonnegative matrix by a matrix in the transport polytope.
+The technique can be readily extended to the Birkhoff polytope.
 ROUND($\boldsymbol{P}$) function for $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$ :
@@
     where $C=\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}$.
-It can be shown that ROUND$(\boldsymbol{P})\in\mathcal{S}_{\varepsilon/s}^{\text{doub}}$ whenever $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$.
-In Algorithm 1, we provide the algorithm for the task.
-In the following section, we focus on solving the intermediate problem (8), after which we move on to the analysis of the Algorithm.
-Problem (P2) motivates the definition of auxiliary variables $\boldsymbol{\gamma}(1),\boldsymbol{\gamma}(2),\dots\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(1),\boldsymbol{Q}(2),\dots$ to aid the algorithm.
-In particular, we will show that we can use our algorithm to get arbitrarily close to $\phi^{\text{opt}}$ in objective value.
-III-A Solving Problem (8)
-Finding $\boldsymbol{\gamma}(t+1)$: Notice that this problem can be solved due to the Assumption A2.
-Finding $\boldsymbol{\tilde{P}}(t+1)$: We will only consider the case when $t$ is even.
-The case when $t$ is odd can be solved similarly.
-Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$.
-To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
-$x_{i}=\begin{cases}\eta Q_{i}(t)\hat{S}_{i,j}(t)&\text{ if }i\in[n],j\in[m]\\ 0&\text{ otherwise}\end{cases}$
-and $\boldsymbol{y}$ is the $j$-th column of $\boldsymbol{\tilde{P}}(t)$. The problem to be solved is
-(P3:) $\underset{\boldsymbol{p}}{\text{min}}-\sum_{i=1}^{s}x_{i}p_{i}+D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{y})$
-[…]
-where $D_{\text{KL}}(\cdot\lVert\cdot)$ in this case is the KL-divergence.
-It should be noted that (P3) has a classic structure that is solved in [11].
-In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$.
-First, assume that $\boldsymbol{z}$ is sorted in the increasing order.
-Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
-$u^{i}_{j}=\begin{cases}\varepsilon&\text{ if }j\leq i\\ \frac{z_{j}}{\sum_{l=i+1}^{s}z_{l}}(1-\varepsilon i)&\text{ if }j>i\\ \end{cases}$
-satisfies, $u^{i}_{j}\geq\varepsilon$ for all $j\in[i+1:s]$.
-Then it can be shown that $\boldsymbol{u}^{i}$ is the solution to (P3).
-Hence, solving (P3) amounts to calculating $\boldsymbol{u}^{i}$ for each $i\in[0:s-1]$ and checking the above condition.
-We first establish the following two lemmas regarding the ROUND operation. A version of the following lemma is proved for the transport polytope in [33]. We adapt the idea for the Birkhoff polytope.
+We establish the relationships between the input and the output matrices of the ROUND operation.
+A version of the following lemma is proved for the transport polytope in [5].
+We adapt the idea for the Birkhoff polytope.
 Lemma 12
-The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$, intermediate matrices $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function in Algorithm 1 satisfy
+The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function satisfy
   1. 1.
@@
   3. 3.
-    $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$.
-  4. 4.
     $\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right)$.
 Proof:
-See Appendix A. ∎
-Next, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
+See Appendix C
+V-B Analysis of the Algorithm
+The rest of the section focuses on proving Theorem 2.
+Similar to Section IV, we consider the interval $T_{0}:T_{0}+T-1]$, where the success probabilities $q_{i,j}$ remain fixed.
+We first notice that Lemma 6, Corollary 1.1, and Lemma 8 hold in the same form for Algorithm 2.
+These results are used in the proof of Theorem 2.
+The intermediate problem (55) admits an explicit solution for $\boldsymbol{\tilde{P}}(t+1)$, as discussed in Appendix D.
+First, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
 Lemma 13
@@
 Combining the two claims, we are done. ∎
-Now, we introduce some useful preliminary lemmas. First, we introduce the push-back lemma regarding minimizing strongly convex functions (See, for example [51]).
-Lemma 4 (Push-back)
-Consider $\mathcal{Q}\in\{\mathcal{S}^{\text{doub}}_{\varepsilon},\mathcal{S}^{\text{row}}_{\varepsilon},\mathcal{S}^{\text{col}}_{\varepsilon},\mathcal{S}\}$, and $\varepsilon\in[0,1/s]$.
-Let $g:\mathbb{R}_{+}^{s\times s}\to\mathbb{R}$ be a convex function.
-Fix $\alpha>0$, and $\boldsymbol{Y}\in\mathcal{Q}$.
-Let
-$\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}\left[g(\boldsymbol{X})+\alpha D(\boldsymbol{X}\lVert\boldsymbol{Y})\right],$
-where $D(\cdot\lVert\cdot)$ is the divergence defined in (15). Then,
-$g(\boldsymbol{X}^{*})+\alpha D(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}),$
-for all $\boldsymbol{Z}\in\mathcal{Q}$.
-We now establish the following two lemmas.
+Now, we first introduce the following lemma that bounds the error introduced by replacing the probabilities $q_{i,j}$ with $\hat{S}_{i,j}(t)$.
 Lemma 14
-We have for $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}$,
-$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert^{2},$ (20)
-where $D(\boldsymbol{X}\lVert\boldsymbol{Y})$ is the divergence defined in (15).
-Proof:
-Notice that
-$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\sum_{i=1}^{s}\frac{1}{2}\left(\sum_{j=1}^{s}|X_{i,j}-Y_{i,j}|\right)^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}$
-where the first inequality follows from the Pinsker’s inequality and the last inequality follows from the fact that for $a_{1},a_{2},\dots,a_{s}\in\mathbb{R}$, $\sum_{i=1}^{s}a_{i}^{2}\geq\left(\sum_{i=1}^{s}a_{i}\right)^{2}/s$.
-∎
-Next, we state a result on the properties of divergence D.
-Lemma 6
-We have $D(\boldsymbol{X}\lVert\boldsymbol{Y})\leq s\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{X}\in\mathcal{S}$, where $\boldsymbol{Y}\in\mathcal{S}_{\varepsilon}$.
-Proof:
-Notice that
-$D(\boldsymbol{X}\lVert\boldsymbol{Y})$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right)\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{1}{\varepsilon}\right)=s\log\left(\frac{1}{\varepsilon}\right),$
-where (a) follows since $X_{i,j}\leq 1$, $Y_{i,j}\geq\varepsilon$, and log is a non-decreasing function. ∎
-Before moving on to the main theorem, we establish a deterministic bound on the queue size $\lVert\boldsymbol{Q}(t)\rVert$ for our Algorithm (Algorithm 1).
-Lemma 7
-We have for all $t\in\mathbb{N}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$, where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1 and $V$ is the utility parameter of Algorithm 1.
-Proof:
-We first prove the following claim.
-Claim: If $Q_{k}(t)>BV$, for $k\in[n]$, then $\gamma_{k}(t+1)=0$.
-Proof:
-Assume the contrary $Q_{k}(t)>BV$, and $\gamma_{k}(t+1)>0$.
-Consider $\boldsymbol{\tilde{\gamma}}$ such that
-$\tilde{\gamma}_{i}=\begin{cases}0&\text{ if }i=k\\ \gamma_{i}(t+1)&\text{ otherwise.}\end{cases}$
-Since, $Q_{k}(t)>BV$, we have
-$[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (21)
-due to the bounded subgradient property (Assumption A1). Hence, notice that
-$\sum_{i=1}^{n}$ $[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})=\gamma_{k}(t+1)[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (22)
-where the last inequality follows due to $\gamma_{k}(t+1)>0$ from assumption and (21).
-The subgradient inequality for the convex function $f:[0,1]^{n}\to\mathbb{R}$ given by, $f(\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}$ gives,
-$\sum_{i=1}^{n}[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})\leq f(\boldsymbol{\gamma}(t+1))-f(\boldsymbol{\tilde{\gamma}})\leq 0,$
-where the last inequality follows from the optimality of $\boldsymbol{\gamma}(t+1)$ for (P3). This is a contradiction with (22). ∎
-Now, we move on to the proof of Lemma 7.
-Fix $i\in[n]$.
-We use induction to prove that $Q_{i}(t)\leq BV+1$ for all $t\in\mathbb{N}$.
-Notice that the result trivially follows for $t=1$.
-Assume $Q_{i}(t)\leq BV+1$.
-We establish $Q_{i}(t+1)\leq BV+1$.
-We consider two cases.
-Case 1: $Q_{i}(t)>BV$: Notice that from the previous claim, we should have $\gamma_{i}(t+1)=0$. Hence, we have from the queuing equation (16),
-$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)\leq BV+1,$
-where (a) follows since $\gamma_{i}(t+1)=0$, and $X_{i}(t)\geq 0$, and the last inequality follows from induction hypothesis.
-Case 2: $Q_{i}(t)\leq BV$: We have from the queuing equation (16),
-$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)+1\leq BV+1,$
-where (a) follows since $\gamma_{i}(t+1)\leq 1$, and $X_{i}(t)\geq 0$, and the last inequality follows from the description of the case.
-Hence, we are done. ∎
-Define the drift $\Delta(t)$ as
-$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
-and the history $\mathcal{H}(t)$ as the sigma algebra generated by
-[…]
-Notice that $\boldsymbol{\tilde{P}}(\tau),\boldsymbol{P}(\tau),\boldsymbol{\gamma}(\tau),\boldsymbol{Q}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable.
-Now, we have the following bound on $\Delta(t)$.
-Lemma 8
-We have for all $t\in\mathbb{N}$,
-$\Delta(t)$ $\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
-Proof:
-Notice that from the queuing equation (16), we have for all $i\in[n]$,
-$Q^{2}_{i}(t+1)$ $\leq\left(Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\right)^{2}$
-$\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t+1)+X^{2}_{i}(t)+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right]$
-$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right],$
-where for the last inequality, we have used $X_{i}(t)\in\{0,1\}$. Summing the above for $i\in[n]$, we have
-$\lVert\boldsymbol{Q}(t+1)\rVert^{2}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right].$
-Taking the expectations conditioned $\mathcal{H}(t)$:
-$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\mathbb{E}\{\gamma_{i}(t+1)|\mathcal{H}(t)\}-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right].$
-Taking expectations, and performing simple algebraic manipulations, we have the result. ∎
-Lemma 9
-We have
-$-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$ $\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
-$\ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
+We have for any $t\in[T_{0},\,T_{0}+T-1]$,
+- $\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$
+$\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
 where $\hat{S}_{i,j}(t)$ is defined in line 6 of Algorithm 2.
@@
 $\sum_{i=1}^{n}\sum_{j=1}^{m}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\}$
-$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
-$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (24)
+$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}$
+$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (60)
 Now, we handle each of the above four terms separately. Notice that
@@
 $\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
-$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{12s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}$
-$\leq_{(a)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}$
-$\leq_{(b)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
-where (a) follows from Lemma 3, and (b) follows from Lemma 5. Next, notice that
-$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$ $=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}$
-$=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
+$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}}{12s\eta}$
+$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}}{3s\eta}$
+$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}}{3\eta},$
+where the second inequality follows from Lemma 13, and the last inequality follows from Lemma 3. Next, notice that
+$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$
+$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
 Now, using the above in (14) we have
@@
 $\leq\frac{nms}{\varepsilon}(BV+1)^{2},$
-where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 3, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Lemma 7. Combining with (9), we are done. ∎
+where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 13, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Corollary 1.1. Combining with (14), we are done. ∎
 Now, we introduce the following lemma that will be useful in deriving the final bounds.
@@
 Proof:
-Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
+Adding $V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
 $\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
@@
 where (a) follows by Lemma 14.
-Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (8) (see Algorithm 1) with Lemma 4, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
+Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (55) (see Algorithm 2) with Lemma 2-1, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
 $-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)$
@@
 $\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
-$\ \ \ \ =-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}P_{i,j}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
-$\ \ \ \ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
+$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}-\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
 Substituting the above in (15), we have
@@
 where for the last inequality, we have used Lemma 4.
-Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Lemma 7, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done.
+Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Corollary 1.1, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done.
 ∎
-Next, we have the following lemma that combines iterates $\boldsymbol{\gamma}(T_{0}+1),\boldsymbol{\gamma}(T_{0}+2)\dots,\boldsymbol{\gamma}(T+T_{0})$, and $\boldsymbol{X}(T_{0}),\boldsymbol{X}(T_{0}+1)$ $\dots,\boldsymbol{X}(T+T_{0}-1)$.
-Lemma 11
-We have for any $T,T_{0}\in\mathbb{N}$,
-$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T}.$
-Proof:
-Notice that
-$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)$
-$=\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right)\right)$
-$\leq_{(a)}\phi\left(\frac{1}{T}\Bigg{(}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}[\boldsymbol{Q}(t+1)-\boldsymbol{Q}(t)]\right)\Bigg{)}$
-$\leq_{(b)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{\boldsymbol{Q}(T+T_{0})}{T}\right)$
-$\leq_{(c)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T},$
-where (a) follows from the entrywise nondecreasing property of $\phi$, and the queuing equation (16), for (b) we use $\boldsymbol{Q}(T_{0})\geq 0$ and the entrywise nondecreasing property of $\phi$, and (c) follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.
-[^2] Assumption A1 ensures that $\phi$ can be extended to an entrywise nondecreasing function that is $\sqrt{nB}$-Lipschitz continuous over the domain $[0,\infty)^{n}$.
-Now, we are ready to establish the performance bound of the algorithm.
-Theorem 1
-For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/s)$, $\eta,V>0$, Algorithm 1 yields
-$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$ $\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}$
-$\ \ \ \ +\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}.$
-In particular
-  1. 1.
-    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
-    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
-  2. 2.
-    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
-    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
+Now, we are ready to establish Theorem 2
 Proof:
@@
 $\leq_{(a)}nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\varepsilon nmsT(BV+1)+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2},$ (64)
-where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Lemma 7.
+where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Corollary 1.1.
 Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain
@@
 $\ \ \ \ +\frac{nB(BV+1)}{T},$
-where (a) follows due to Lemma 11, and (b) follows due to Lemma 7. Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{\text{opt}}$. Using Jensen’s inequality, we are done. ∎
-IV Discussion
-IV-A Adaptiveness
-A careful inspection of the proof of Theorem 1 shows that the result holds even if the success probabilities $q_{i,j}$ changed before time $T_{0}$, as long as they remained constant during $[T_{0}:T_{0}+T-1]$.
-Hence, the adaptiveness is satisfied.
-IV-B Enforcing user fairness
-Consider the case when $n\geq m$ (the number of users is at least the number of channels).
-Assume we fix a $\theta\in(0,1/s]$ such that each user is required to transmit at least $\theta$ fraction of the time on average on each channel and no user can transmit more than $(1-(n-m)\theta)$ fraction of the time.
-This enables faster adaptation to the new optimality point.
-To see this, notice that for (P1), we require the additional constraints of
-$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y_{i,j}(t)\}\geq\theta.$
-for all $i\in[n]$ and $j\in[m]$ and
-$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}\{Y_{i,j}(t)\}\leq 1-(n-m)\theta,$
-for all $i\in[n]$.
-This will simply transform the constraint $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$ of (P2) to $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\theta}$.
-Setting $\varepsilon=\theta$ in Algorithm 1, this will allow us to use $\boldsymbol{P}^{*}$ directly in (1) in Theorem 1, instead of $(1-\varepsilon s)\boldsymbol{P}^{*}+\varepsilon\boldsymbol{1}$, which gives
-$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
-$\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\theta V}(BV+1)^{2}+\frac{\theta nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\theta}\right)+\frac{n(BV+1)^{2}}{2VT}$
-$\ \ \ \ +\frac{nB(BV+1)}{T}.$
-Now, since we assume $\theta$ to be a constant, using $\eta=\Theta(1/T)$, and $V=\Theta(\sqrt{T})$, we have
-$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}=O\left(\frac{1}{\sqrt{T}}\right)$
-IV-C Case $m=1$
-This case can be approached using the work on bandits with vector rewards and concave utility functions [26].
-However, the algorithms proposed are not adaptive.
-It turns out that the algorithm for the general multi-channel case can be simplified for the single-channel case preserving the adaptiveness.
-In addition, the simplified algorithm has faster convergence.
-The idea of the algorithm is to find a distribution $\boldsymbol{p}(t)$ over the users in the $(t-1)$-th iteration, after which we sample the user to be assigned on the $t$-th iteration from $\boldsymbol{p}(t)$.
-Similar to the multi-channel case, we use auxiliary variables $\boldsymbol{\gamma}(t)\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(t)\in\mathbb{R}^{n}_{+}$ for all $t\in\mathbb{N}$.
-For notational convenience, for each $t\in\mathbb{N}$ and $i\in[n]$, we will use $q_{i},S_{i}(t)$ and $Y_{i}(t)$ instead of $q_{i,1},S_{i,1}(t)$ and $Y_{i,1}(t)$, respectively.
-The algorithm uses three parameters $\varepsilon\in(0,1/n)$, $V>0$, and $\eta>0$.
-See Algorithm 2 for details of the implementation.
-1 Initialize $\boldsymbol{p}(1)\in\Delta_{n,\varepsilon}$, $\boldsymbol{\gamma}(1)\in[0,1]^{n}$, and the virtual queues $\boldsymbol{Q}\in[0,BV+1]^{n}$.
-2 for each iteration $t\in[1:T]$ do
-3       Sample $a_{t}\sim\boldsymbol{p}(t)$ and set $\boldsymbol{Y}(t)=\boldsymbol{e}_{a_{t}}$ where $\boldsymbol{e}_{i}$ is the unit vector with $i$-th entry being 1.
-4       Receive feedback $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$.
-5       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using, $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[1:n]$.
-6       Find, $\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1)$ by solving the problem,
-$(\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1))=\arg\min_{\boldsymbol{p}\in\Delta_{n,\varepsilon},\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})$
-$+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big{]},$ (29)
-where $D_{\text{KL}}$ is the KL-divergence.
-7       Update the virtual queues,
-$\boldsymbol{Q}(t+1)=[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)]_{+},$ (30)
-where $\boldsymbol{X}(t)=\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
-8
-Algorithm 2 Single Channel Adpative MAC
-Most of the results required for the error analysis are borrowed from the multi-channel case.
-We only prove the results that are unique to this case.
-We first focus on solving the intermediate problem (6).
-Notice that the problem can be separated into the problem of finding optimal $\boldsymbol{\gamma}(t+1)$ and $\boldsymbol{p}(t+1)$.
-Finding $\boldsymbol{\gamma}(t+1)$: This is the same problem as in the multi-channel case.
-Finding $\boldsymbol{p}(t+1)$: This is (P3) with $\boldsymbol{x}=\eta\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)$, and $\boldsymbol{y}=\boldsymbol{p}(t)$.
-Similar to Lemma 5, we have Pinsker’s inequality for KL divergence.
-Lemma 12 (Pinsker’s inequality)
-For $\boldsymbol{x},\boldsymbol{y}\in\Delta_{n}$, we have that,
-$D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert_{1}^{2}\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert^{2},$
-We also have the following lemma similar to Lemma 6.
-Lemma 13
-We have $D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\leq\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{x}\in\Delta_{n}$ where $\boldsymbol{y}\in\Delta_{n,\varepsilon}$.
-Notice that since the queue updates and the problem to solve $\boldsymbol{\gamma}(t+1)$ are the same as in the multi-channel case, we have the same queue bound.
-Lemma 14
-We have that for all $t\in\{1,2,,\dots\}$ and $i\in\{1,2\dots,n\}$, $Q_{i}(t)\leq BV+1$.
-Define the drift $\Delta(t)$ as,
+where (a) follows due to Lemma 11, and (b) follows due to Corollary 1.1.
+Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{*}\geq\phi^{\text{opt}}$.
+Combining with Jensen’s inequality, we are done.
+VI Special Cases
+In this section, we introduce several interesting cases of the general setting.
+The first is an adaptive algorithm for the single-channel case ($m=1$).
+This algorithm has both closed form solutions for iterations and fast convergence.
+Then we introduce a UCB-based non-adaptive algorithm with efficient implementation.
+Both algorithms will be used for comparisons.
+Finally, we consider the single channel ($m=1$) special case with $\phi(\boldsymbol{x})=\min\{x_{1},x_{2},\dots,x_{n}\}$.
+In this setting, a simple mechanism achieves convergence to the optimal value while adapting to channel variations, without requiring explicit estimation.
+VI-A Single Channel Algorithm
+The single channel case simplifies due to the absence of matching constraints.
+In particular, in each iteration, we assign a user to the channel, and the assigned user in the $t$-th time slot is sampled from $\boldsymbol{p}(t)\in\Delta^{n}$.
+For this section, we use the notation $S_{i}(t)=S_{i,1}(t)$, $Y_{i}(t)=Y_{i,1}(t)$.
+Algorithm 3 summarizes the steps.
+1Initialize $\boldsymbol{p}(1)=\frac{1}{n}\boldsymbol{1}$ and the virtual queues $\boldsymbol{Q}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$.
+2 for each time slot $t\in\mathbb{N}$ do
+3   Sample $l_{t}\in[n]$ from $\boldsymbol{p}(t)$, set $\boldsymbol{Y}(t)$ as the $l_{t}$-th unit vector, and receive feedback.
+4   Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[n]$
+5   Find
+$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big[-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big].$
+$\boldsymbol{p}(t+1)=\arg\min_{\boldsymbol{p}\in\Delta^{n}_{\varepsilon}}\Big[-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big],$ (65)
+6   Update $\boldsymbol{Q}(t+1)=\left[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right]_{+}$.
+Algorithm 3 Single-Channel Adaptive MAC (Parameters $V>0$, $\varepsilon\in(0,1/n])$, and $\eta>0$)
+We first state the performance bound of the algorithm.
+Theorem 3
+Consider $T,T_{0}\in\mathbb{N}$ and suppose that the success probabilities $q_{i,j}$ remain fixed on the interval $T_{0},\,T_{0}+T-1]$.
+For any parameters $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
+$\phi^{\text{opt}}$ $-\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$
+$\leq\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
+In particular, using $\eta=\Theta(\sqrt{\log(T)}/T)$, $\varepsilon=\Theta(\sqrt{\log(T)}/\sqrt{T})<1$, and $V=\Theta(T^{1/2})$, we have
+$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\sqrt{\frac{\log(T)}{T}}\right).$
+VI-A1 Solving Problem (5)
+Notice that finding $\boldsymbol{\gamma}(t+1)$ is the same problem as in the multi-channel case.
+In addition, the problem to solve for $\boldsymbol{p}(t+1)$ reduces to the problem (P3) discussed in Appendix D.
+VI-A2 Deterministic Problem
+Consider the problem
+$\underset{\boldsymbol{p},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (66)
+$\text{s.t. }\boldsymbol{p}\in\Delta_{n}$ (67)
+$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (68)
+$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (69)
+Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is at least $\phi^{\text{opt}}$.
+For this section, let $\phi^{*}$ denote the optimal objective value of the above problem.
+Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote an optimal solution of the above problem.
+If $q_{1},q_{2},\dots,q_{n}$ were known, the policy of sampling a user independently in each time slot using the distribution $\boldsymbol{p}^{*}$ achieves an objective value of $\phi^{*}$ for (P1).
+The subsequent analysis shows that even when $q_{1},q_{2},\dots,q_{n}$ are unknown, we can achieve an objective value arbitrarily close to $\phi^{*}$ for (P1), thereby showing $\phi^{*}=\phi^{\text{opt}}$.
+VI-A3 Analysis of the Algorithm
+Most of the results required for the error analysis are borrowed from the multi-channel case. In particular, notice that the deterministic queue bound Corollary 1.1 holds in this case.
+Define the drift $\Delta(t)$ as
 $\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}.$ (70)
@@
 Lemma 17
-We have that,
--- $\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
-$\ \ \ \ \leq\frac{\eta n}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$
-where $\hat{S}_{i}(t)$ is defined in line 5 of Algorithm 2.
+For any $t\in[T_{0}:T+T_{0}-1]$, we have
+- $\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}$
+$\leq-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (71)
+where $\hat{S}_{i}(t)$ is defined in line 4 of Algorithm 3.
 Proof:
-Notice that,
-$\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\}$
-$=\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)[p_{i}(t+1)-p_{i}(t)]\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
-$\leq_{(a)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{2\eta}\mathbb{E}\left\{\lVert\boldsymbol{p}(t+1)-\boldsymbol{p}(t)\rVert^{2}\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
-$\leq_{(b)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$ (32)
-where (a) follows from $\frac{1}{2\eta}\lVert\boldsymbol{a}\rVert^{2}+\frac{\eta}{2}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{n}a_{i}b_{i}$ for $n$-dimensional vectors $\boldsymbol{a,b}$, (b) follows from Lemma 12.
-Define the history $\mathcal{H}(t)$ similar to multi-channel case (see, (23)).
-Notice that $\boldsymbol{p}(\tau),\boldsymbol{Q}(\tau),\boldsymbol{\gamma}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable.
-Now, we handle the terms of (16) separately.
-Notice that,
-$\mathbb{E}$$\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}^{2}(t)\right\}$
-$=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}\hat{S}_{a_{t}}(t)\right\}\leq_{(a)}\frac{1}{\varepsilon}\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}(t)\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\right\}$
-$=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=_{(b)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}\mathbb{E}\left\{S_{i}(t)Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}$
-$=_{(c)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}q_{i}\mathbb{E}\left\{Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(d)}\frac{1}{\varepsilon}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
-$\leq\frac{n}{\varepsilon}(BV+1)^{2}$
-where $a_{t}$ is defined in line 3 of Algorithm 2, (a) follows since $S_{a_{t}}(t)\leq 1$ and $p_{a_{t}}(t)\geq\varepsilon$, (b) follows since $\boldsymbol{Q}(t)$ and $\boldsymbol{p}(t)$ are $\mathcal{H}(t)$-measurable, (c) follows since $\boldsymbol{S}(t)$ is independent of $\boldsymbol{Y}(t)$ and $\mathcal{H}(t)$, (d) follows since $q_{i}\leq 1$, and the last inequality follows due to Lemma 14.
-Also,
-$\sum_{i=1}^{n}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}=\mathbb{E}\left\{Q_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}p_{a_{t}}(t)\right\}=\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
-Combining everything and substituting in (16), we are done. ∎
+Using Lemma 5, we have
+$\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
+$\leq_{(a)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(\sum_{i=1}^{n}{p}_{i}(t)e^{\eta Q_{i}(t)\hat{S}_{i}(t)}\right)\Bigg]$
+$\leq_{(b)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(\sum_{i=1}^{n}{p}_{i}(t)\left[1+\eta Q_{i}(t)\hat{S}_{i}(t)+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\right]\right)\Bigg]$
+$=\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(1+\sum_{i=1}^{n}\eta{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}p_{i}(t)Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\right)\Bigg]$
+$\leq_{(c)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}\eta{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}p_{i}(t)Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\Bigg]$
+$=_{(d)}\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}S_{i}(t)Y_{i}(t)\hat{S}_{i}(t)$
+$=_{(e)}\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}\sum_{i=1}^{n}\hat{S}_{i}(t)$ (72)
+where (a) follows from Lemma 5, (b) follows combining Corollary 1.1 and $\hat{S}_{i}(t)\leq 1/\varepsilon$ with $e^{x}\leq 1+x+(e^{c}-c-1)x^{2}/c^{2}$ for $x\in[0,c]$ for any $c>0$, (c) follows from $\ln(1+x)\leq x$, (d) follows from the definition of $\hat{S}_{i}(t)$, and the last inequality follows since $S_{i}(t),Y_{i}(t)\in\{0,1\}$.
+Taking the expected value conditioned on $\mathcal{H}(t)$, followed by taking unconditional expectations, we have
+$\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
+$\leq\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}\sum_{i=1}^{n}q_{i}$
+$\leq\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (73)
+where the last inequality follows since $q_{i}\in[0,1]$. Rearranging, we are done. ∎
 Now, we introduce the following lemma.
 Lemma 18
-We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
+We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$,
 $VT\phi(\boldsymbol{\gamma})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
-$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
+$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (74)
 Proof:
-Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 15, we have that,
+Adding $V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 16, we have that,
 $\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
-$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
-$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{\eta n}{2\varepsilon}(BV+1)^{2}$
-$\ \ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$ (33)
-where (a) follows by Lemma 16. Now notice that combining the decision (6) with Lemma 4, we have that for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
-$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
-$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))-\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1)).$
-Taking expectations of the above, we have that,
-$-V$ $\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}$
-$\ \ \ \ \ -\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
-$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
-$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
-$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}p_{i}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
-$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}.$
-Substituting the above in (17), we have that,
+$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$ (75)
+Using Lemma 17 on the above, we have
 $\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
-$\leq n+\frac{\eta n}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
-$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
-Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, we have that,
+$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
+$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (76)
+Notice that from the optimality of $(\boldsymbol{\gamma}(t+1),\boldsymbol{p}(t+1))$ in (5), we have that for any $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$ and $\boldsymbol{\gamma}\in[0,1]^{n}$
+$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
+$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}$
+$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}p_{i}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}$
+$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$
+$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}-\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$ (77)
+where the inequality follows from Lemma 2-2. Using the above in (18), we have for any $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$ and $\boldsymbol{\gamma}\in[0,1]^{n}$
+$\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq n+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}$
+$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$ (78)
+Now, we add the above for $t\in[T_{0}:T_{0}+T-1]$ to get
 $\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
-$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\right\}$
-$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\right\}$
-$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right),$
-where for the last inequality, we have used Lemma 13.
-Rearranging the above and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done.
+$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\}$
+$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\}$
+$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\}$
+$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\}$
+$\leq_{(a)}nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$ (79)
+where (a) follows by $D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{s})\leq\log(1/\varepsilon)$ for $\boldsymbol{p}\in\Delta^{n}$ and $\boldsymbol{s}\in\Delta^{n}_{\varepsilon}$ (Proof follows the same steps as Lemma 4).
+Rearranging and using $\lVert Q(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ (follows from Corollary (1.1)), we are done.
 ∎
-Now, we are ready to establish the performance bound of the algorithm.
-Theorem 2
-For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
-$\phi^{\text{opt}}$ $-\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$
-$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
-In particular
-  1. 1.
-    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
-    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
-  2. 2.
-    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
-    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
+Now, we are ready to establish Theorem 3.
 Proof:
-Consider the problem
-$\underset{\boldsymbol{p},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (34)
-[…]
-$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (36)
-$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (37)
-Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is $\phi^{\text{opt}}$.
-Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote the solution of the above problem.
-Substituting $\boldsymbol{p}^{*}(1-\varepsilon n)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 17 we have that,
-$VT\phi^{\text{opt}}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
-$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i}(1-\varepsilon n)-\varepsilon q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
-$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}nq_{i}p^{*}_{i}\mathbb{E}\{Q_{i}(t)\}$
+Substituting $(1-\varepsilon n)\boldsymbol{p}^{*}+\varepsilon,\boldsymbol{\gamma}^{*}$ in Lemma 18, where $\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*}$ is the optimal solution of Problem (66)-(69), we have that,
+$VT\phi^{*}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
+$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-p^{*}_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon n\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\}$
 $\ \ \ \ +\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
-$\leq_{(a)}nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\varepsilon n^{2}T(BV+1)+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (38)
-where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (34)-(37)), and Lemma 18.
-Now, we use the Jensen’s inequality to obtain,
-$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
-$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
-Notice that Lemma 11 holds since the queue updates are the same as in the multi-channel case. Combining above with Lemma 11, we have that,
+$\leq_{(a)}nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\varepsilon n^{2}(BV+1)T+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (80)
+where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (66)-(69)), and Corollary (1.1).
+Dividing both sides by $VT$ and using Jensen’s inequality, we have
+$\phi^{*}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
+$\leq\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$ (81)
+Combining above with Lemma 11, we have that,
 $\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
-$\leq_{(a)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
-$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
-$\leq_{(b)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
-where (a) follows due to Lemma 11, and (b) follows due to Lemma 18. Now, using Jensen’s inequality, we are done. ∎
-Comparing the error bounds of Theorem 2 and Theorem 1 (with $m=1$), it can be seen that the bound of Theorem 2 is strictly better.
-In particular, the second term of the error bound has a $\Theta(n)$ dependence on $n$ in Theorem 2, whereas the corresponding value is $\Theta(n^{3})$ in Theorem 1.
-V UCB Algorithm
+$\leq_{(a)}\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
+$\leq_{(b)}\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
+where (a) follows due to Lemma 11, and (b) follows due to Corollary (1.1). Now, using Jensen’s inequality, we are done. ∎
+VI-B Non-Adaptive MAC
 In this section, we present a UCB-based algorithm for the link selection problem in MAC (Algorithm 4), after which we move on to the analysis.
-Recall that $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ is the user-channel assignment during time slot $t$.
 For $t\in[s]$, we set $\boldsymbol{Y}(t)$ arbitrarily such that each user-channel pair is explored exactly once during the interval $[s]$.
-For each $t>s$, we compute a permutation matrix $\boldsymbol{M}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1.
-The assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ for the $t>s$ is obtained by discarding the last $s-m$ columns or the last $s-n$ rows of $\boldsymbol{M}(t)$ depending on whether $n$ or $m$ is larger.
+For each $t>s$, we compute permutation matrix $\boldsymbol{Y}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1.
 The algorithm uses parameters $\delta_{s},\delta_{s+1},\dots$ that satisfy $\delta_{t}>0$ for all $t\in\{s,s+1,\dots\}$.
@@
 $\text{UCB}_{i,j}(t)=\hat{S}_{i,j}(t)+f_{i,j}(t),\ \forall t\in\{s,s+1,\dots\}$ (82)
-1 For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
+1For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
 2
@@
 4 for each time slot $t\in\{s+1,s+2,\dots,\}$ do
-5       Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{M}(t)$ by solving
+5   Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{Y}(t)$ by solving
 $\boldsymbol{\gamma}(t)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{N}Q_{i}(t)\gamma_{i},$ (83)
-$\boldsymbol{M}(t)=\arg\max_{\boldsymbol{M}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
-where $\text{UCB}_{i,j}(t-1)$ is defined in (V).
-6       Use the assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$, where $Y_{i,j}(t)=M_{i,j}(t)$ for $(i,j)\in[n]\times[m]$ and receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
+$\boldsymbol{Y}(t)=\arg\max_{\boldsymbol{Y}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
+where $\text{UCB}_{i,j}(t-1)$ is defined in (VI-B).
+6   Receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
 7   Update the queues
@@
 Algorithm 4 UCB MAC
+We first state the performance bound of Algorithm 4.
+Theorem 4
+Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
+$\phi^{\text{opt}}$ $-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
+$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
+$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (85)
+where $f_{i,j}(t)$ is defined in (VI-B) and $\phi^{\text{max}}=\max_{\boldsymbol{\gamma}\in[0,1]^{n}}\phi(\boldsymbol{\gamma})$. In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
+  1. 1.
+    For all $V>0$,
+    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
+    $\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{V(T-s)}$
+    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
+  2. 2.
+    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
+    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right),$
+    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
+VI-B1 Solving Problem (83)
 We first focus on solving the inner problem (83).
 Notice that the problem to solve to obtain $\boldsymbol{\gamma}(t)$ is the same as for Algorithm 1.
-The problem to solve to obtain $\boldsymbol{M}(t)$ has a classic max-weight structure (hence, $\boldsymbol{M}(t)$ is a permutation matrix).
+The problem to solve to obtain $\boldsymbol{Y}(t)$ has a classic max-weight structure (hence, $\boldsymbol{Y}(t)$ is a permutation matrix).
 Hence, we can use the Hungarian algorithm [32] to obtain the solution.
-Now, we move on to the analysis of the Algorithm 3.
+VI-B2 Analysis of the Algorithm
+Now, we focus on proving Theorem 4.
 Since the queue updates and the problem to solve to obtain $\boldsymbol{\gamma}(t)$ are the same as in Algorithm 1, we have the same deterministic queue bound.
@@
 We have for all $t\in\{s+1,s+2,\dots\}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$
-Proof:
-This follows repeating the argument of Lemma 7. ∎
 Define the collection of good events $\Omega_{s+1},\Omega_{s+2},\dots$ as
-[…]
-where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (V). We have the following lemma.
+$\Omega_{t}=\Bigg\{$ $\text{UCB}_{i,j}(t-1)-2f_{i,j}(t-1)\leq q_{i,j}\leq\text{UCB}_{i,j}(t-1),\forall i\in[n],j\in[m]\Bigg\},$ (86)
+where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (VI-B). We have the following lemma.
 Lemma 20
@@
 Proof:
-The proof follows directly by applying the Hoeffding inequality with union bound. See [48] for more details. ∎
+The proof follows directly by applying the Hoeffding inequality with union bound. See [lattimore_szepesvári_2020] for more details. ∎
 Define the drift $\Delta_{t}$
@@
 and history $\mathcal{H}(t)$ as
-[…]
+$\mathcal{H}(t)=\{$ $\boldsymbol{Y}(1),\dots,\boldsymbol{Y}(t-1),\boldsymbol{Y}(1)\odot\boldsymbol{S}(1),\dots,\boldsymbol{Y}(t-1)\odot\boldsymbol{S}(t-1)\},$ (87)
 for each $t\in\{s+1,s+2,\dots\}$. Notice that $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable.
@@
 Proof:
-Proof is similar to the proof of Lemma 8 and is given in Appendix B. ∎
+Notice that from the queuing equation (84), we have
+$Q^{2}_{i}(t+1)$ $\leq(Q_{i}(t)+\gamma_{i}(t)-X_{i}(t))^{2}\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t)+X^{2}_{i}(t)+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
+$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
+for all $i\in[n]$. Summing the above for $i\in[n]$, we have
+$\lVert\boldsymbol{Q}(t+1)\rVert^{2}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)].$
+Taking the expectations conditioned on the history $\mathcal{H}(t)$ defined in (87), we have
+$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\mathbb{E}\{\gamma_{i}(t)|\mathcal{H}(t)\}-\mathbb{E}\{X_{i}(t)|\mathcal{H}(t)\}]$
+$=\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right],$
+where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result. ∎
 Fix a time horizon $T\in\{s+1,s+2,\dots\}$. Then we have the following result.
@@
 $\ \ \ \ +\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t},$
-where $f_{i,j}(t)$ is defined in (V).
+where $f_{i,j}(t)$ is defined in (VI-B).
 Proof:
@@
 $\leq_{(a)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg|\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)P_{i,j}^{*}\Bigg|\Omega_{t}\right\}$
-$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}\leq-V\phi^{\text{opt}}$ (45)
-where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{M}(t)$ for intermediate problem (40) in Algorithm 3 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (42); the last inequality follows from (11).
+$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg|\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg|\Omega_{t}\right\}$
+$\leq_{(c)}-V\phi^{*}\leq-V\phi^{\text{opt}}$ (89)
+where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{Y}(t)$ for the intermediate problem (83) in Algorithm 4 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (86); (c) follows from the constraint (15); the last inequality follows from Lemma 1.
 Adding the inequalities (22) and (22) and rearranging, we are done.
@@
 $\ \ \ \ +2(2V\phi^{\text{max}}+nBV+n)nm\sum_{t=s}^{T-1}\delta_{t}.$
-Dividing both sides by $V(T-s)$, and using Jensen’s inequality, we are done. ∎
+Dividing both sides by V(Ts)V(T-s), and using Jensen’s inequality, we are done. ∎
 Now we have the following lemma that combines $\boldsymbol{\gamma}(s+1),\boldsymbol{\gamma}(s+2),\dots,\boldsymbol{\gamma}(T)$ with $\boldsymbol{X}(s+1),\boldsymbol{X}(s+2),\dots,\boldsymbol{X}(T)$.
@@
 The proof uses the same argument as Lemma 11. ∎
-Now, we have the following theorem that establishes the performance bound of Algorithm 3.
-Theorem 3
-Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
-$\phi^{\text{opt}}$ $-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
-$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
-$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (47)
-where $f_{i,j}(t)$ is defined in (V). In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
-  1. 1.
-    For all $V>0$,
-    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
-    $\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{V(T-s)}$
-    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
-  2. 2.
-    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
-    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right),$
-    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
-  3. 3.
-    Fix $\epsilon>0$. Using $V=\Theta(1/\epsilon)$, we have
-    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}(\epsilon),$
-    and $\mathcal{O}$ hides the dependence on all parameters but $\epsilon$.
+Now, we are ready to prove Theorem 4.
 Proof:
@@
 $\ \ \ \ +\frac{nB(BV+1)}{T-s}.$ (92)
-Combining (3) and (3) and using the Jensen’s inequality, we have (3).
+Combining (VI-B2) and (VI-B2) and using the Jensen’s inequality, we have (4).
 To prove part 1, first fix $i\in[n]$ and $j\in[m]$. When $\delta_{t-1}=1/(t-1)$ for all $t\in[s+1:T]$, we have
@@
 $\ \ \ \ +\frac{2nB(BV+1)}{T}+\frac{(\sqrt{n}B+1)s}{T},$
-which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$.
-For part 3, notice that from part 1,
-$\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)\leq\frac{n}{V}.$
-Using $V=\Theta(1/\epsilon)$, we are done. ∎
-VI Discussion
-VI-A Distributed Implementation
-Consider a system where the $n$ users take decisions in the absence of a centralized controller.
-In each time slot, each user receives as feedback whether each channel was successfully accessed or not (this can be, for instance, via sensing each channel).
-Notice that $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$ are $\mathcal{H}(t)$-measurable, where $\mathcal{H}(t)$ is defined in (43).
-Hence, during time slot $t$, if each user has access to the full history $\mathcal{H}(t)$, they can independently implement Algorithm 3 to find $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$.
-From the feedback, each user can compute $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$ and hence the full history $\mathcal{H}(t+1)$.
-Hence, starting from $\mathcal{H}(s+1)$, each user can independently implement Algorithm 3.
-Notice that this is not possible in Algorithm 1, since the channel assignments are not $\mathcal{H}(t)$-measurable.
+which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$. ∎
+VI-C Single-Channel Case with Minimum Utility
+Consider the single-channel case $m=1$ with utility function $\phi(\boldsymbol{x})=\min\{x_{1},\dots,x_{n}\}$.
+Observe that if $q_{i}=0$ for some $i\in[n]$, then $\phi\left(\mathbb{E}\{\boldsymbol{\bar{X}}(T)\}\right)=0$ under any assignment policy.
+Consequently, $\phi^{\text{opt}}=0$.
+Therefore, without loss of generality, we assume that $q_{i}>0$ for all $i\in[n]$.
+If the values $q_{1},\dots,q_{n}$ are known, an optimal policy for (P1) is to sample user $i$ with probability $p_{i}^{*}=\frac{1/q_{i}}{\sum_{k=1}^{n}(1/q_{k})}$.
+Under this policy, $\mathbb{E}[X_{i}(t)]=q_{i}p_{i}^{*}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}$ for all $i\in[n]$.
+Hence, we have $\phi^{\mathrm{opt}}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}$.
+Now consider the following procedure:
+  • Sample a user uniformly at random.
+  • Continue to serve the user until a successful transmission.
+  • Resample a user uniformly at random and repeat.
+This mechanism does not require knowledge of the $q_{i}$.
+Fix a user $i\in[n]$.
+Let $(T_{j},R_{j})_{j\geq 1}$ denote the renewal process where, $T_{j}$ is the number of slots required until the $j$-th sampled user achieves a success, $R_{j}=1$ if the $j$-th sampled user is $i$, and $R_{j}=0$ otherwise.
+Note that the sequence ${(T_{j},R_{j})}_{j\geq 1}$ is independent and identically distributed.
+Since users are sampled uniformly at random, and conditional on sampling user $k$, the transmission duration is geometrically distributed with mean $1/q_{k}$, we have $\mathbb{E}\{R_{1}\}=\frac{1}{n}$ and $\mathbb{E}[T_{1}]=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{q_{k}}$.
+By the renewal reward theorem, with probability 1,
+$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}X_{i}(t)=\frac{\mathbb{E}[R_{1}]}{\mathbb{E}[T_{1}]}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}=\phi^{\text{opt}}.$ (96)
+Thus every user achieves throughput $\phi^{\text{opt}}$, and therefore
+$\lim_{T\to\infty}\min\{\bar{X}_{1}(T),\dots,\bar{X}_{n}(T)\}=\phi^{\mathrm{opt}}$ (97)
+with probability 1. Hence, the mechanism converges to the optimal value without requiring knowledge of $q_{1},\dots,q_{n}$.
+Suppose now that the values $q_{i}$ change at some time.
+Because geometric service times are memoryless, the residual service time at the change point has the same distribution as a fresh geometric random variable with the new parameter.
+Therefore, from the change point onward, the system evolves as a delayed renewal process under the new parameters, and the renewal argument still applies.
+Hence, the process is adaptive.
 VII Simulations
-For simulations, we consider four scenarios with $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+x_{i})$, $n=5,m=3$, and $T=10^{5}$.
-In each scenario, the values of $q_{i,j}$ change mid-way through the simulation (at $T=5\times 10^{4}$).
-In Scenario 1 (Figure 1-Left-Top), certain user-channel links are turned off at $T=5\times 10^{4}$.
-In Scenario 2 (Figure 1-Right-Top), certain user-channel links that were turned off before, turn on at $T=5\times 10^{4}$.
-In Scenarios 3 (Figure 1-Left-Bottom) and 4 (Figure 1-Right-Bottom), the values of $q_{i,j}$ for certain user-channel links change at $T=5\times 10^{4}$.
-In each scenario, the algorithms are not told about the change.
-For each figure, for each algorithm, we plot $\phi\left(\frac{1}{t}\sum_{\tau=1}^{t}\boldsymbol{X}(\tau)\right)$ for $t\leq T/2$ and $\phi\left(\frac{1}{t-T/2}\sum_{\tau=T/2+1}^{t}\boldsymbol{X}(\tau)\right)$ for $t>T/2$ vs $t$.
-We also plot the optimal objective value of (P2), for reference.
-Note that in the first half of the simulation, both algorithms converge to the optimal solution in all scenarios, although the convergence of Algorithm 3 is faster.
-However, to show the importance of Algorithm 1, we observe the second half of the simulation.
-Although Algorithm 3 achieves superior performance in the first half of the simulation, it does not perform well in the second half except in Scenario 2.
-In contrast Algorithm 1 achieves similar performance in both halves in all scenarios.
-This is due to the adaptiveness of Algorithm 1.
-Figure 1: Top-Left: Scenario 1. Top-Right: Scenario 2. Bottom-Left: Scenario 3. Bottom-Right: Scenario 4.
+We consider eight simulation scenarios.
+All experiments were conducted on a virtual machine equipped with an Intel Xeon 2.20 GHz CPU (2 virtual CPUs).
+The algorithms were implemented in Python.
+In the first four scenarios, we use the smooth utility function $\phi(\mathbf{x})=\sum_{i=1}^{n}\log(1+b_{i}x_{i})$, where $b_{i}>0$ are fixed constants.
+In scenarios 5 and 6, we consider the nonsmooth utility function $\phi(\mathbf{x})=a\sum_{i=1}^{n}x_{i}+b\min(x_{1},x_{2},\dots,x_{n})$, where $a,b>0$, are fixed constants.
+In the last two scenarios, we use the utility function $b\min(x_{1},x_{2},\dots,x_{n})$, where $a,b>0$, is a fixed constant.
+We run each algorithm for $T=10^{6}$ time slots.
+The channel success probabilities are changed once at time $T_{0}=5\times 10^{5}$, in order to evaluate the adaptiveness of the algorithms.
+The algorithms are neither restarted nor informed of this change.
+To examine performance before and after the change point, we plot the evolution of the objective applied to empirical averages of the observed vectors $\boldsymbol{X}(t)$.
+We compute
+$\Phi(t)=\begin{cases}\phi\!\left(\frac{1}{t}\sum_{\tau=1}^{t}\mathbf{X}(\tau)\right),&\text{ if }t<T_{0}\\ \phi\!\left(\frac{1}{t-T_{0}+1}\sum_{\tau=T_{0}}^{t}\mathbf{X}(\tau)\right)&\text{ if }t\geq T_{0}\end{cases}$ (98)
+We then plot $\Phi(t)$ as a function of $t$.
+This representation allows us to separately examine the convergence behavior before the change and the transient adaptation after the change.
+In each plot, we also include the optimal utility value obtained by solving Problem (P1) using the corresponding channel success probabilities.
+Figure 1 denotes the plots.
+In each plot, the y-axis is zoomed in to clearly highlight the performance differences among the algorithms.
+If the curve corresponding to a particular algorithm is not visible in the second half of the plot, it indicates that the algorithm has not adapted sufficiently to reach the new optimal value.
+(a) Scenario 1
+(b) Scenario 2
+(c) Scenario 3
+(d) Scenario 4
+(e) Scenario 5
+(f) Scenario 6
+(g) Scenario 7
+(h) Scenario 8
+Figure 1: Evolution of the objective $\Phi(t)$ defined in (98). A vertical dashed line marks $T_{0}=5\times 10^{5}$, the point at which the success probabilities change.
+In Scenarios 1 and 2, we consider a multi-channel setting with $n=4$ and $m=3$.
+In Scenario 1, the success probabilities $q_{i,j}$ in the two halves of the simulation are chosen randomly.
+In Scenario 2, we interchange these $q_{i,j}$ values.
+Here, we compare the Adaptive MAC (Algorithm 1), Adaptive MAC.CF (Algorithm 4), and UCB-MAC (Algorithm 4) algorithms.
+Notice that the adaptive algorithms have decent performance in the first half of the simulation.
+Although UCB-MAC has faster convergence in the first half, it does not adapt to changes in success probabilities in the second half.
+An interesting observation is that, despite the stronger worst-case guarantee of the Adaptive MAC algorithm, the Adaptive MAC.CF algorithm appears to outperform it in simulation.
+In Scenarios 3 and 4, we consider a single-channel setting with $n=4$.
+We assume that some links may become unavailable without the controller’s knowledge; in that case, the corresponding $q_{i,j}$ is set to zero.
+When a link is available, its success probability is $q_{i,j}=0.4$.
+Within each half of the simulation horizon, the availability status of each link remains fixed.
+Here, for the adaptive algorithm, we use the Single-Channel Adaptive MAC (Algorithm 3).
+We also compare with the Algorithms introduced in [2] (Algorithm 1 of [2], and Frank-Wolfe-based Algorithm 3 of [2]).
+In this case, the adaptation capability of the adaptive MAC algorithm can be clearly seen.
+In fact, in Scenario 3, the UCB-MAC algorithm performs extremely poorly in the second half of the simulation, resulting in an objective value that drops close to zero.
+For Scenario 5 and Scenario 6, we again consider a single channel scenario.
+Recall that the utility function in this case is $\phi(\boldsymbol{x})=a\sum_{i=1}^{n}x_{i}+b\min(x_{1},x_{2},\dots,x_{n})$.
+We compare Single-Channel Adaptive MAC, UCB-MAC, and Algorithm 1 of [2].
+Notice that even for the nonsmooth utility, the adaptive algorithm adapts fast in the second half of the simulation.
+In Scenarios 7 and 8, we consider the single-channel setting with utility $\phi(\boldsymbol{x})=b\min(x_{1},x_{2},\dots,x_{n})$.
+We also plot the simple renewal mechanism described in Section VI-C.
+In these scenarios, UCB-MAC also adapts in the second half of the simulation.
+Interestingly, we observe that UCB-MAC does not rely on explicit UCB-based estimation in this case.
+Instead, it effectively implements a mechanism that closely resembles the renewal mechanism.
+In Table II, we report the average runtime per time slot of the algorithms under different scenarios.
+We observe that Adaptive MAC.CF has faster iterations than Adaptive MAC.
+However, the UCB-based non-adaptive algorithm outperforms both adaptive algorithms in terms of runtime.
+In the single-channel scenario, the adaptive algorithm achieves runtimes comparable to those of the UCB MAC algorithm.
+Finally, the UCB-based non-adaptive algorithm outperforms the algorithms proposed in [2] in terms of empirical runtime complexity.
+TABLE II: Average runtime per time slot(in milliseconds) of each algorithm under each scenario.
+For the Adaptive MAC algorithm, we also report (in parenthesis) the average number of Sinkhorn iterations per main iteration.
+Algorithm Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7 Scenario 8
+Adaptive 3.878 3.849 - - - - - -
+MAC (12.982) (12.054) - - - - - -
+Adaptive 2.655 2.433 - - - - - -
+MAC.CF
+UCB MAC 0.207 0.211 0.157 0.171 0.0682 0.07 0.0651 0.0655
+Single
+Channel - - 0.242 0.27 0.0729 0.0744 0.0706 0.0715
+MAC
+BwCR
+Algorithm 1 - - 9.224 9.556 0.0691 0.0773 0.0672 0.0655
+[2]
+BwCR
+Algorithm 3 - - 0.17 0.171 - - - -
+[2]
 VIII Conclusions
-This paper focused on the problem of designing algorithms for automatic link selection in multi-channel multiple access with link failures.
-In particular, we solved a network utility maximization problem with matching constraints and bandit feedback on link failures.
-We considered two algorithms, where the first algorithm has slower convergence and is adaptive, and the second algorithm has faster convergence and is not adaptive.
-Appendix A Proof of Lemma 2
+This paper focuses on the problem of designing algorithms for automatic link selection in multi-channel multiple access systems with transmission failures.
+In particular, we solve a network utility maximization problem with matching constraints and bandit feedback on transmission outcomes.
+We propose two adaptive algorithms: the first achieves faster convergence, while the second admits a more efficient implementation.
+We also study several special cases with additional structural properties.
+Extending the framework to adversarial settings and incorporating more general combinatorial constraints are directions for future work.
+Appendix A SINKHORN function
+We introduce SINKHORN proceure on positive $s\times s$ matrices.
+[…]
+Initialize $\boldsymbol{R}^{1}=\boldsymbol{R}$.
+For each iteration $\tau\in\{1,2\dots,\}$
+  1. 1.
+    Row normalization: If $\tau$ is even set $\boldsymbol{R}^{\tau+1}$ as
+    $R_{i,j}^{\tau+1}=\frac{R_{i,j}^{\tau}}{\sum_{k=1}^{s}R_{i,k}^{\tau}}$ (99)
+  2. 2.
+    Column normalization: Else set $\boldsymbol{R}(\tau+1)$ as
+    $R_{i,j}^{\tau+1}=\frac{R_{i,j}^{\tau}}{\sum_{k=1}^{s}R_{k,j}^{\tau}}$ (100)
+Given $\boldsymbol{X}\in\mathbb{R}^{s\times s}$, and $\boldsymbol{Q}\in(0,\infty)^{s\times s}$, SINKHORN studied in [24] is an iterative process to solve problems of the type
+$\min_{\boldsymbol{P}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}P_{i,j}+D(\boldsymbol{P}\lVert\boldsymbol{Q})$ (101)
+We provide the following lemma adapted from [24]
+Lemma 24
+Given $\epsilon>0$, running SINKHORN($\boldsymbol{R}$), where $\boldsymbol{R}$ is defined by $R_{i,j}=Q_{i,j}\exp(-X_{i,j})$ for sufficiently many iterations (say $U$ iterations) generates a matrix $\boldsymbol{R}^{U}$ such that
+$\exp(-\epsilon)\leq\frac{R^{U}_{i,j}}{P^{*}_{i,j}}\leq\exp(\epsilon)$ (102)
+where $\boldsymbol{P}^{*}$ is the solution of Problem (101).
+Appendix B Finding $\boldsymbol{\hat{P}}(t+1)$ in Problem (26)
+Finding $\boldsymbol{\hat{P}}(t+1)$ can be summarized under the following steps.
+  1. 1.
+    Define matrix $\boldsymbol{R}\in(0,\infty)^{s\times s}$ by $R_{i,j}=P_{i,j}(t)\exp(\eta Q_{i}(t)\hat{S}_{i,j}(t))$
+  2. 2.
+    Run SINKHORN($\boldsymbol{R}$) for $U$ iteration to get $\boldsymbol{R}^{U}$ such that
+    $e^{-\theta/(3s)}\leq\frac{R^{U}_{i,j}}{\tilde{P}_{i,j}(t+1)}\leq e^{\theta/(3s)}$ (103)
+    This is possible due to Lemma 24. See Appendix A for details.
+  3. 3.
+    Set $\boldsymbol{\hat{P}}(t+1)=\text{ROUND}(\boldsymbol{R}^{U})$. See Section V-A for details.
+We have the following lemma that establishes the required property of $\boldsymbol{\hat{P}}(t+1)$.
+Lemma 25
+We have that $\boldsymbol{\hat{P}}(t+1)$ found above is a $(\theta/s)$-approximation of $\boldsymbol{\tilde{P}}(t+1)$
+Proof:
+First, notice that for each $i\in[s]$
+$\sum_{k=1}^{s}R_{i,k}^{U}\leq\sum_{k=1}^{s}\tilde{P}_{i,j}(t+1)e^{\theta/(3s)}=e^{\theta/(3s)},$ (104)
+where the inequality follows from (103), and the equality follows since $\boldsymbol{\tilde{P}}(t+1)$ is doubly stochastic by definition.
+Similarly, for each $j\in[s]$, we have $\sum_{k=1}^{s}R_{k,j}^{U}\leq e^{\theta/(3s)}$.
+First, from Lemma 12, we have that $\boldsymbol{\hat{P}}(t+1)\in\mathcal{S}^{\text{doub}}$. Also, from the description of the ROUND function in Section V-A, notice that
+$\hat{P}_{i,j}(t+1)$ $\geq\frac{R_{i,j}^{U}}{(\sum_{k=1}^{s}R_{i,k}^{U})(\sum_{k=1}^{s}R_{k,j}^{U})}\geq_{(a)}\frac{R_{i,j}^{U}}{e^{2\theta/(3s)}}\geq\frac{\tilde{P}_{i,j}(t+1)}{e^{\theta/s}}$ (105)
+where (a) follows from the previous two inequalities, and (b) follows from (103). Hence, we are done. ∎
+Appendix C Proof of Lemma 12
 Before proving this, notice that due to the normalization steps 1 and 2 of the ROUND function, we have $\boldsymbol{1}\geq\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}$, and $\boldsymbol{1}\geq(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}$, where the inequalities are taken entrywise.
@@
 So we are done.
-3) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$.
+3) We first prove the following claim
+Claim: We have that $\lvert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
+Proof:
+First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$.
 Then we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0$.
@@
 Notice that
-$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
-$=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
+$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
 $=_{(a)}\frac{\boldsymbol{1}^{\top}\left(\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\right)\boldsymbol{1}}{\boldsymbol{1}^{\top}\left(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\right)}$
@@
 where (a) follows since $\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\geq 0$, and $\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}\geq 0$, as a result of which $(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\geq 0$.
-4) Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
-$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s$
-$=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
+∎
+Now we prove part 3. Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
+$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
 $=\sum_{i\in\mathcal{U}}\left(\sum_{j=1}^{s}P_{i,j}-1\right)+\sum_{i\in\mathcal{U}^{c}}\left(1-\sum_{j=1}^{s}P_{i,j}\right)+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
@@
 $\geq_{(c)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left[\sum_{i=1}^{s}P_{i,j}-1\right]_{+}\right)$
-$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}P_{i,j}-1\right\rvert\right)$
+[…]
 $\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$ (106)
@@
 Now, to complete the proof notice that
-$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
-$=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
+$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
 $\leq_{(b)}\lVert\boldsymbol{P}\rVert_{1}-2\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+s\leq_{(c)}\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$
 $\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right),$ (108)
-where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to part 3, and (c) follows from (54).
-Appendix B Proof of Lemma 20
-Notice that from the queuing equation (41), we have
-$Q^{2}_{i}(t+1)$ $\leq(Q_{i}(t)+\gamma_{i}(t)-X_{i}(t))^{2}\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t)+X^{2}_{i}(t)+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
-$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
-for all $i\in[n]$. Summing the above for $i\in[n]$, we have
-$\lVert\boldsymbol{Q}(t+1)\rVert^{2}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)].$
-Taking the expectations conditioned on the history $\mathcal{H}(t)$ defined in (43), we have
-$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\mathbb{E}\{\gamma_{i}(t)|\mathcal{H}(t)\}-\mathbb{E}\{X_{i}(t)|\mathcal{H}(t)\}]$
-$=\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right],$
-where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result.
+where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to the claim, and (c) follows from (107).
+Appendix D Finding $\boldsymbol{\tilde{P}}(t+1)$ in Problem (55)
+We only consider the case where $t$ is even.
+The case where $t$ is odd can be solved similarly.
+Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$.
+To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
+$x_{i}=\begin{cases}\eta Q_{i}(t)\hat{S}_{i,j}(t)&\text{ if }i\in[n],j\in[m]\\ 0&\text{ otherwise}\end{cases}$
+and $\boldsymbol{y}$ is the $j$-th column of $\boldsymbol{\tilde{P}}(t)$. The problem to be solved is
+(P3:) $\underset{\boldsymbol{p}}{\text{min}}-\sum_{i=1}^{s}x_{i}p_{i}+D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{y})$
+$\text{s.t. }\boldsymbol{p}\in\Delta^{s}_{\varepsilon},$
+It should be noted that (P3) has a classic structure that is solved in [27].
+In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$.
+First, assume that $\boldsymbol{z}$ is sorted in the increasing order.
+Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
+[…]
+satisfies, $u^{i}_{j}\geq\varepsilon$ for all $j\in[i+1:s]$.
+Then it can be shown that $\boldsymbol{u}^{i}$ is the solution to (P3).
+Hence, solving (P3) amounts to calculating $\boldsymbol{u}^{i}$ for each $i\in[0:s-1]$ and checking the above condition.